### 30 Mar 2026 v2

# Wasserstein Propagation for Reverse Diffusion under Weak Log-Concavity: Exploiting Metric Mismatch via One-Switch Routing

Abstract
Existing analyses of reverse diffusion typically propagate sampling error in the Euclidean geometry underlying $W_{2}$ throughout the reverse trajectory. Under weak log-concavity, this can be suboptimal: Gaussian smoothing may create contraction first at large separations, while short-scale Euclidean dissipativity is still absent. We show that exploiting this metric mismatch can yield strictly sharper end-to-end $W_{2}$ bounds than direct full-horizon Euclidean propagation on mismatch windows. Our analysis derives an explicit radial lower profile for the learned reverse drift, whose far-field and near-field limits quantify the contraction reserve and the residual Euclidean load, respectively. This profile determines admissible switch times and leads to a one-switch routing theorem: reflection coupling damps initialization mismatch, pre-switch score forcing, and pre-switch discretization in an adapted concave transport metric; a single $p$-moment interpolation converts the damped switch-time discrepancy back to $W_{2}$; and synchronous coupling propagates the remaining error over the late Euclidean window. Under $L^{2}$ score-error control, a one-sided monotonicity condition on the score error, and standard well-posedness and coupling assumptions, we obtain explicit non-asymptotic end-to-end $W_{2}$ guarantees, a scalar switch-selection objective, and a conversion exponent $\theta_{p}=(p-2)/(2(p-1))$ that cannot be improved uniformly within the affine-tail concave class under the same $p$-moment switch assumption. For a fixed switch, the routed and direct Euclidean bounds share the same late-window term, so any strict improvement is entirely an early-window effect.
1 Introduction
Score-based diffusion models are now a central framework for generative modeling, building on early diffusion and score-based formulations (Sohl-Dickstein et al., 2015; Song and Ermon, 2019; Ho et al., 2020; Song et al., 2020). At a conceptual level, sampling is a reverse-time transport problem. The smoothed intermediate laws determine its geometry, score approximation perturbs the drift, and numerical samplers discretize the dynamics. An end-to-end theory of diffusion sampling should therefore identify the geometry that actually governs error propagation along reverse time and turn it into sharp non-asymptotic bounds.
A growing theoretical literature provides convergence guarantees for diffusion samplers in several discrepancy measures, including KL divergence, total variation, and Wasserstein distances (Conforti et al., 2025; Lee et al., 2023; Gao et al., 2025; Gao and Zhu, 2024; Gentiloni-Silveri and Ocello, 2025). We focus on the $2$-Wasserstein distance $W_{2}$, in which initialization mismatch, score approximation, and discretization admit a common coupling-based treatment. In globally regular regimes, discrepancy can often be propagated directly in $W_{2}$ along the entire reverse trajectory, without changing metric across phases (Gao et al., 2025; Gao and Zhu, 2024). Once global regularity is lost, however, this Euclidean route need not remain sharp. The natural question is: what transport geometry actually governs the propagation of sampling error along the reverse trajectory, and how should end-to-end $W_{2}$ bounds reflect that geometry rather than a uniform Euclidean worst case?
We study this question under weak log-concavity, where the target law remains globally confining but may be locally nonconvex, in the spirit of recent Ornstein–Uhlenbeck regularization analyses beyond global log-concavity (Gentiloni-Silveri and Ocello, 2025). In this regime, Gaussian smoothing need not produce Euclidean contractivity uniformly across scales. Instead, the radial lower profile we derive may become contractive first at large separations while remaining obstructed near the origin. A direct full-horizon Euclidean propagation is then bottlenecked by the short-scale obstruction even though contraction is already available farther out. The effective transport geometry is therefore scale-dependent, and treating it as uniformly Euclidean can make the resulting end-to-end $W_{2}$ bound genuinely conservative.
On an early window, discrepancy is propagated in a concave transport metric adapted to the radial profile. At an admissible switch point $s_{0}$ in the forward smoothing variable, the damped discrepancy is returned once to $W_{2}$, and the remaining late window is propagated in Euclidean geometry. For a fixed admissible switch, the routed and direct Euclidean bounds share the same late-window term, so any advantage is entirely an early-window effect; on mismatch windows, the routed bound can be strictly sharper.
Our contributions.
Our contributions are fourfold.
  1. (i)
    We derive an explicit radial lower profile for the learned reverse drift under weak log-concavity. This profile separates far-field contraction from near-field obstruction and yields a transparent admissibility criterion for switch points.
  2. (ii)
    We develop the early-window contraction mechanism behind the routed analysis. Reflection coupling reduces the separation process of two copies of the learned reverse SDE to a one-dimensional radial diffusion, from which we construct an adapted concave switch metric $W_{\varphi_{s_{0}}}$ that propagates initialization mismatch, pre-switch score forcing, and early discretization up to the switch.
  3. (iii)
    We prove an end-to-end one-switch $W_{2}$ routing theorem. The theorem localizes metric mismatch at a single interface, returns the pre-switch discrepancy from $W_{\varphi_{s_{0}}}$ to $W_{2}$ only once under a $p$-moment budget, induces a scalar switch-selection objective over admissible grid-aligned switches, and yields the exponent $\theta_{p}=(p-2)/(2(p-1))$, which cannot be improved uniformly within the affine-tail concave class under the same switch-moment assumption.
  4. (iv)
    We give an exact comparison with the standard direct full-horizon Euclidean route. For any fixed admissible switch, the routed and direct bounds share the same late-window term, so the comparison reduces entirely to the early window. This identifies mismatch windows on which the routed bound can be strictly sharper.
1.1 Technique overview
The proof is organized around a single geometric fact. Beyond globally regular regimes, reverse diffusion is controlled not by one uniform Euclidean dissipativity constant but by a scale-dependent radial contractivity profile of the learned reverse drift. Under weak log-concavity, Gaussian smoothing can make this profile contractive first at large separations even while the Euclidean one-sided monotonicity bound required for direct $W_{2}$ propagation remains non-contractive near the origin. The first usable contraction therefore appears in a transport geometry different from the Euclidean geometry in which the final error is measured.
Let $p_{s}$ denote the smoothed data law at noise level $s$, and write s=Tts=T-t for forward time along the reverse trajectory. In the score-based SDE formulation (Song et al., 2020), the ideal and learned reverse dynamics are
$\mathop{}\!\mathrm{d}\widetilde{X}_{t}$ $=b_{t}(\widetilde{X}_{t})\,\mathop{}\!\mathrm{d}t+g(T-t)\,\mathop{}\!\mathrm{d}\overline{B}_{t},$ (1)
$\mathop{}\!\mathrm{d}\widehat{Y}_{t}$ $=\widehat{b}_{t}(\widehat{Y}_{t})\,\mathop{}\!\mathrm{d}t+g(T-t)\,\mathop{}\!\mathrm{d}\overline{B}_{t},$
where $b_{t}$ is the ideal reverse drift and $\widehat{b}_{t}$ is the learned reverse drift. For a vector field $u$, define its radial drift profile by
$\kappa_{u}(r):=\inf_{\|x-y\|=r}-\frac{\langle u(x)-u(y),x-y\rangle}{\|x-y\|^{2}},\qquad r>0.$
Under Assumptions 1 and 3—the latter being a one-sided monotonicity condition on the score error rather than a global Lipschitz assumption—the radial profile dictionary in Section 3 yields
$\kappa_{\widehat{b}_{t}}(r)\;\geq\;g^{2}(s)\Big(\alpha_{s}-\frac{1}{r}f_{M_{s}}(r)-\ell(s)\Big)-f(s),\qquad s=T-t,$ (2)
with explicit coefficients determined by smoothing and score-error control. We denote the right-hand side of (2) by $\underline{\kappa}_{s}(r)$.
A single lower envelope controls both geometric regimes used later:
$\lim_{r\downarrow 0}\underline{\kappa}_{s}(r)=-\mathfrak{b}(s),\qquad\lim_{r\to\infty}\underline{\kappa}_{s}(r)=\mathfrak{m}(s).$
Thus $\mathfrak{b}(s)$ and $\mathfrak{m}(s)$ are not separate bookkeeping quantities; they are the near-field and far-field ends of the same radial lower profile. The near-field endpoint records the residual Euclidean obstruction, while the far-field endpoint records the contraction reserve created by Gaussian smoothing.
From synchronous to reflection coupling.
A direct $W_{2}$ propagation uses synchronous coupling. Because the two copies are driven by the same Brownian motion, the noise cancels in the distance dynamics, so a Euclidean Grönwall closure requires a global one-sided bound
$\langle x-y,\widehat{b}_{t}(x)-\widehat{b}_{t}(y)\rangle\leq\beta(t)\|x-y\|^{2},\qquad x,y\in\mathbb{R}^{d}.$ (3)
In globally regular regimes, such a closure is typically obtained from a global one-sided Lipschitz control of the learned score or reverse drift (Gao et al., 2025; Gao and Zhu, 2024). The restriction is that (3) must hold uniformly over all radii. Synchronous coupling is therefore governed by the worst part of the radial profile $r\mapsto\kappa_{\widehat{b}_{t}}(r)$, which in the early weakly log-concave regime is precisely the near field. The direct Euclidean route thus fails not because large-scale contraction is absent, but because it cannot exploit that contraction before every radius is already Euclideanly dissipative.
To use the full scale-dependent profile, we instead employ reflection coupling (Eberle, 2016; Eberle et al., 2019). Fix an admissible switch $s_{0}$, and compress the early geometry into a switch-level lower envelope $\underline{\kappa}_{s_{0}}$, negative near the origin and positive beyond a threshold $R_{\mathrm{sw}}(s_{0})$. If $Y_{t}$ and $\bar{Y}_{t}$ are reflected copies of the learned reverse SDE and $r_{t}:=\|Y_{t}-\bar{Y}_{t}\|$, then, up to coupling time, the multidimensional dynamics reduce schematically to
$\mathop{}\!\mathrm{d}r_{t}\;\lesssim\;-\underline{\kappa}_{s_{0}}(r_{t})\,r_{t}\,\mathop{}\!\mathrm{d}t+2g(T-t)\,\mathop{}\!\mathrm{d}W_{t}.$ (4)
Unlike synchronous coupling, the radial dynamics retains a Brownian term and therefore sees the entire switch-level profile rather than only its worst short-scale Euclidean value. This is exactly why a concave radial cost becomes effective: applying Itô’s formula to $\varphi(r_{t})$ produces the second-order term $2g(T-t)^{2}\varphi^{\prime\prime}(r_{t})$, which can offset the near-field load.
The switch metric as an adapted radial Lyapunov gauge.
The switch metric $\varphi_{s_{0}}$ is therefore not an auxiliary softer cost chosen after the fact. It is the radial Lyapunov gauge naturally associated with the reflected generator. Writing $\underline{g}(s_{0}):=\inf_{u\in[s_{0},T]}g(u),$ the relevant radial operator has the schematic form
$\mathcal{L}_{s_{0}}\varphi(r)=2\underline{g}(s_{0})^{2}\varphi^{\prime\prime}(r)-\underline{\kappa}_{s_{0}}(r)\,r\,\varphi^{\prime}(r).$
We construct $\varphi_{s_{0}}$ as an explicit concave subsolution of
$2\underline{g}(s_{0})^{2}\varphi^{\prime\prime}(r)-\underline{\kappa}_{s_{0}}(r)\,r\,\varphi^{\prime}(r)\leq-c(s_{0})\varphi(r).$ (5)
This is the conceptual core of the construction: the metric is adapted to the reflected radial generator before it is used as a transport cost.
The resulting switch metric is
$\varphi_{s_{0}}(r)=\begin{cases}\displaystyle\int_{0}^{r}e^{-\lambda(s_{0})u^{2}}\,\mathop{}\!\mathrm{d}u,&0\leq r\leq R_{\mathrm{sw}}(s_{0}),\\[8.00003pt] \displaystyle\varphi_{s_{0}}(R_{\mathrm{sw}}(s_{0}))+\mathfrak{a}(s_{0})\bigl(r-R_{\mathrm{sw}}(s_{0})\bigr),&r\geq R_{\mathrm{sw}}(s_{0}),\end{cases}$ (6)
for switch-dependent parameters $\lambda(s_{0})$ and $\mathfrak{a}(s_{0})$. Its form mirrors the two ends of the profile. On $0\leq r\leq R_{\mathrm{sw}}(s_{0})$, where $\underline{\kappa}_{s_{0}}$ may still be negative, the curvature term $2\underline{g}(s_{0})^{2}\varphi_{s_{0}}^{\prime\prime}$ supplied by the reflected radial noise compensates the near-field obstruction. On $r\geq R_{\mathrm{sw}}(s_{0})$, where the profile carries positive radial margin, no curvature is needed; the metric is continued affinely to preserve sensitivity to large separations and hence retain the far-field reserve.
Applying Itô’s formula to $\varphi_{s_{0}}(r_{t})$, together with (5), yields the early semigroup contraction
$W_{\varphi_{s_{0}}}\!\bigl(\mu P_{u,v},\nu P_{u,v}\bigr)\leq e^{-c(s_{0})(v-u)}\,W_{\varphi_{s_{0}}}(\mu,\nu),\qquad 0\leq u\leq v\leq t_{s}:=T-s_{0},$ (7)
where $P_{u,v}$ denotes the learned reverse semigroup. Thus the far-field reserve is converted into genuine early-window transport contraction, but in an adapted concave metric rather than directly in $W_{2}$.
This contraction compresses the entire early window into a single damped switch-time quantity through the same one-step defect interface later used in the Euclidean phase. Since $\varphi_{s_{0}}(r)\leq r$, one has
$W_{\varphi_{s_{0}}}(\mu,\nu)\leq W_{1}(\mu,\nu)\leq W_{2}(\mu,\nu),$
so each one-step $W_{2}$ defect also controls the corresponding $W_{\varphi_{s_{0}}}$ defect.
The one-time return to $W_{2}$ and what it costs.
The switch should not be read as the time at which the geometry suddenly becomes Euclidean. Its role is different: it is the point at which an already damped discrepancy is returned once to the Euclidean metric in which the final error is measured. This return is possible because the switch metric has a global linear lower bound,
$\varphi_{s_{0}}(r)\geq\mathfrak{a}(s_{0})\,r,\qquad r\geq 0,$
and hence
$W_{1}(\mu,\nu)\leq\mathfrak{a}(s_{0})^{-1}W_{\varphi_{s_{0}}}(\mu,\nu).$
Combined with a $p$-moment budget at switch time, this yields the interpolation
$W_{2}^{2}\!\bigl(\operatorname{Law}(Z_{t_{s}}),p_{s_{0}}\bigr)\;\lesssim\;\frac{\rho}{\mathfrak{a}(s_{0})}\,\Delta_{\varphi}^{\mathrm{sw}}(s_{0})+\frac{\overline{\mathcal{M}}_{p}^{\mathrm{sw}}(s_{0})}{\rho^{p-2}},$
and optimizing in $\rho$ gives
$W_{2}\!\bigl(\operatorname{Law}(Z_{t_{s}}),p_{s_{0}}\bigr)\;\lesssim\;C_{p}^{\mathrm{sw}}(s_{0})\,\bigl(\Delta_{\varphi}^{\mathrm{sw}}(s_{0})\bigr)^{\theta_{p}},\qquad\theta_{p}=\frac{p-2}{2(p-1)}.$
This is the only place where the exponent $\theta_{p}$ enters. It measures the price of recovering quadratic Euclidean transport from a switch metric that is deliberately softer than $W_{2}$.
The conversion is also where the central tradeoff becomes explicit. A more concave early metric is better at absorbing the near-field obstruction under the reflected radial dynamics, but within the affine-tail concave class this comes with weaker quadratic sensitivity at large distances and hence a more expensive return to $W_{2}$. The sharpness result shows that this loss is structural rather than a proof artifact: the same non-Euclidean softness that makes early damping possible also prevents a free recovery of Euclidean quadratic transport.
Late-window Euclidean propagation and optimal switching.
After the switch, the remaining interval is propagated in the target metric $W_{2}$, for which synchronous coupling is again the natural tool. This yields
$W_{2}\!\bigl(\operatorname{Law}(Z_{t_{N}}),p_{0}\bigr)\;\lesssim\;\Delta_{\mathrm{late}}(s_{0})+e^{\Gamma(s_{0})}\,C_{p}^{\mathrm{sw}}(s_{0})\,\bigl(\Delta_{\varphi}^{\mathrm{sw}}(s_{0})\bigr)^{\theta_{p}}.$
Admissibility and optimization therefore play different roles. Admissibility is a geometric existence statement: it certifies that a genuinely non-Euclidean contractive regime is available on the early window. Optimal switching is a routing statement: it decides when the already compressed discrepancy should be re-exposed to $W_{2}$, balancing the remaining early damping, the one-time conversion cost, and the residual late Euclidean load.
Organization.
Section 2 positions the paper relative to existing diffusion-model convergence theory. Section 3 introduces the reverse-diffusion setting, the structural assumptions, and the radial geometry of the learned drift. Section 4 states the main theorem and explains its meaning. Section 5 returns to the geometric message and the open directions.
2 Related Work
Our work lies at the intersection of reverse-time diffusion representations (Anderson, 1982; Haussmann and Pardoux, 1986), Wasserstein convergence theory for score-based generative models, analyses of reverse-time geometry beyond global log-concavity, and reflection-coupling contraction methods for nonconvex diffusions.
$W_{2}$ theory under regular and changing geometry.
A broad theoretical literature studies score-based diffusion sampling under minimal assumptions, polynomial-complexity regimes, manifold-supported data, and KL/path-space metrics (Azangulov et al., 2024; Conforti et al., 2025; Chen et al., 2022, 2023a; Lee et al., 2023). Within Wasserstein control, strongly log-concave and globally regular settings admit direct Euclidean propagation arguments for reverse SDEs and PF-ODEs (Bruno et al., 2025; Gao et al., 2025; Gao and Zhu, 2024). More recent $W_{2}$ analyses move beyond global log-concavity, including Gaussian-tail, semiconvex, weak-semiconvex, and weakly log-concave regimes (Wang and Wang, 2024; Conforti, 2024; Bruno and Sabanis, 2025; Gentiloni-Silveri and Ocello, 2025; Kremling et al., 2025). The closest comparisons from our viewpoint emphasize different manifestations of time-varying geometry. Gentiloni-Silveri and Ocello (2025) shows that weak log-concavity can produce alternating contractive and non-contractive phases along reverse time; Bruno and Sabanis (2025) studies semiconvex settings with time-dependent Euclidean $W_{2}$ control; and Kremling et al. (2025) proves weak-log-concave guarantees for PF-ODEs under Euclidean/Lipschitz control. Our result is complementary to these works. The additional point we emphasize is not only when contractivity appears, but also in which metric it first becomes exploitable: a reverse trajectory may already contract in a softer radial transport geometry even while direct Euclidean closure is still blocked by the near field. This metric-mismatch principle is what leads to our one-switch route and is also what allows the routed bound to be sharper than direct full-horizon Euclidean propagation on mismatch windows.
Reflection coupling and concave transport metrics.
Methodologically, we build on the reflection-coupling program for diffusions with nonconvex geometry, from the classical multidimensional constructions of Lindvall and Rogers (1986); Chen and Li (1989) and the coupling overview of Kallenberg (1993) to the concave-distance contraction framework of Eberle (2016); Eberle et al. (2017, 2019). Because our pre-switch analysis compares learned and exact reverse dynamics with different drifts, the sticky-coupling analysis of Eberle and Zimmer (2016) is also methodologically relevant. We use the Gaussian-smoothed radial profile of the learned reverse drift to determine an early-window concave transport metric, propagate initialization mismatch, score forcing, and discretization errors on that window, and then perform a single return to $W_{2}$. Our distinctive contribution is therefore the reverse-diffusion interface and one-switch routing theorem, not reflection coupling in isolation.
Sampler interfaces.
A separate line analyzes deterministic, stochastic, and higher-order diffusion samplers (Chen et al., 2023b; Li et al., 2024; Beyler and Bach, 2025; Yu and Yu, 2025; Pfarr et al., 2026), together with sharper complexity and stability guarantees (Benton et al., 2023; Chen et al., 2026; Strasman et al., 2026). Our theorem is modular with respect to those discretization results. Any one-step or accumulated $W_{2}$ defect estimate enters only through the interface term $\xi_{k}$, after which the routing theorem determines how much of that sampler error is damped on the early radial window and how much is exposed to the late Euclidean window. Thus our contribution is not a new sampler analysis per se, but a propagation theorem that can incorporate sharper sampler bounds without changing the geometric argument.
3 Reverse-diffusion interface and time-dependent radial geometry
Under weak log-concavity, Gaussian smoothing can stabilize large separations before short scales satisfy a Euclidean one-sided dissipativity bound. We encode this scale dependence through a radial profile of the learned reverse drift, whose two endpoints will later determine the residual Euclidean load and the far-field contraction reserve.
Notation.
We use the Euclidean inner product and norm throughout, write $\mathcal{P}_{p}(\mathbb{R}^{d})$ for laws with finite $p$-th moment, and denote by $\Pi(\mu,\nu)$ the set of couplings of probability measures $\mu,\nu$ on $\mathbb{R}^{d}$. We write $W_{2}$ for the $2$-Wasserstein distance and parametrize reverse time by t:=Tst:=T-s. Thus larger $s$ corresponds to earlier reverse time, whereas smaller $s$ corresponds to later reverse time.
3.1 Forward smoothing and reverse interface
Let $p_{0}(\mathop{}\!\mathrm{d}x)\propto e^{-V_{0}(x)}\,\mathop{}\!\mathrm{d}x$, where $V_{0}\in C^{1}(\mathbb{R}^{d})$ and $\int_{\mathbb{R}^{d}}e^{-V_{0}(x)}\,\mathop{}\!\mathrm{d}x<\infty$. Given Borel measurable $f,g:[0,T]\to\mathbb{R}_{\geq 0}$ with $\int_{0}^{T}f(u)\,\mathop{}\!\mathrm{d}u<\infty$, $\int_{0}^{T}g^{2}(u)\,\mathop{}\!\mathrm{d}u<\infty$, and $g(s)>0$ for every $s>0$, consider the forward diffusion
$\mathop{}\!\mathrm{d}X_{s}=-f(s)X_{s}\,\mathop{}\!\mathrm{d}s+g(s)\,\mathop{}\!\mathrm{d}B_{s},\qquad X_{0}\sim p_{0}.$ (8)
Writing
$a(s):=\exp\!\Big(-\int_{0}^{s}f(u)\,\mathop{}\!\mathrm{d}u\Big),\qquad\sigma^{2}(s):=\int_{0}^{s}\Big(\frac{a(s)}{a(u)}\,g(u)\Big)^{2}\,\mathop{}\!\mathrm{d}u,$ (9)
its marginal law is
$p_{s}=(S_{a(s)}p_{0})*\mathcal{N}(0,\sigma^{2}(s)I_{d}),\qquad s>0,$ (10)
where $S_{a}\mu$ is the pushforward of $\mu$ under $x\mapsto ax$. Thus forward time combines deterministic linear shrinkage with Gaussian convolution. The corresponding reverse drifts are
$b_{t}(x)=f(T-t)x+g^{2}(T-t)\nabla\log p_{T-t}(x),\qquad\widehat{b}_{t}(x)=f(T-t)x+g^{2}(T-t)s_{\theta}(x,T-t),$ (11)
and the ideal and learned reverse SDEs are Anderson (1982); Haussmann and Pardoux (1986)
$\mathop{}\!\mathrm{d}\widetilde{X}_{t}$ $=b_{t}(\widetilde{X}_{t})\,\mathop{}\!\mathrm{d}t+g(T-t)\,\mathop{}\!\mathrm{d}\overline{B}_{t},\qquad\widetilde{X}_{0}\sim p_{T},$ (12)
$\mathop{}\!\mathrm{d}\widehat{Y}_{t}$ $=\widehat{b}_{t}(\widehat{Y}_{t})\,\mathop{}\!\mathrm{d}t+g(T-t)\,\mathop{}\!\mathrm{d}\overline{B}_{t},\qquad\widehat{Y}_{0}\sim\widehat{p}_{T}.$ (13)
With
$e_{s}(x):=s_{\theta}(x,s)-\nabla\log p_{s}(x),\qquad s\in(0,T],$
the drift mismatch is exactly
$\widehat{b}_{t}-b_{t}=g^{2}(T-t)e_{T-t}.$ (14)
Whenever the learned reverse SDE (13) is well posed on an interval $u,v]\subseteq[0,T]$, we write $P_{u,v}$ for its transition kernel. Background facts on time reversal, regularity, and well posedness are collected in Appendix A.
Definition 1 (Strong well posedness on an interval).
We say that the learned reverse SDE is strongly well posed on an interval $u,v]\subseteq[0,T]$ if for every $\mathcal{F}_{u}$-measurable initial condition $\xi$ with $\operatorname{Law}(\xi)\in\mathcal{P}_{1}(\mathbb{R}^{d})$, there exists a pathwise unique strong solution on $[u,v]$.
3.2 Structural assumptions
Assumption 1 (Weak log-concavity Gentiloni-Silveri and Ocello (2025)).
Let $\alpha,M>0$, and define $f_{M}(r):=2\sqrt{M}\,\tanh\!\Big(\frac{1}{2}\sqrt{M}\,r\Big).$ Assume that for every $r>0$,
$\inf_{\|x-y\|=r}\frac{\langle\nabla V_{0}(x)-\nabla V_{0}(y),x-y\rangle}{\|x-y\|^{2}}\geq\alpha-\frac{1}{r}f_{M}(r).$
Assumption 2 (Score forcing in $L^{2}(p_{s})$).
Assume there exists a measurable function $\varepsilon:(0,T]\to[0,\infty)$ such that
$\mathop{\mathbb{E}}\nolimits_{X\sim p_{s}}\|e_{s}(X)\|^{2}\leq\varepsilon^{2}(s),\qquad s\in(0,T],$
and $s\mapsto g^{2}(s)\varepsilon(s)$ is integrable on $(0,T]$.
Assumption 3 (One-sided monotonicity control of the score error).
Assume Assumption 2 holds, and that there exists a measurable function $\ell:(0,T]\to\mathbb{R}$ such that
$\langle e_{s}(x)-e_{s}(y),x-y\rangle\leq\ell(s)\|x-y\|^{2},\qquad x,y\in\mathbb{R}^{d}.$
Remark 1 (On the score-error assumptions).
Assumption 2 controls averaged score forcing, whereas Assumption 3 controls the pairwise geometric component of the score error that enters distance evolution through $\langle e_{s}(x)-e_{s}(y),x-y\rangle$. Thus Assumption 3 is a one-sided monotonicity condition rather than a full global regularity requirement. Global Lipschitz or symmetrized Jacobian bounds are stronger sufficient conditions implying it, but are not intrinsic to the theorem. The two assumptions are complementary rather than interchangeable. Concrete sufficient conditions, together with a comparison against global Lipschitz-type hypotheses, are given in Appendix G.
3.3 A radial profile dictionary and admissible switches
What drives the proof is not a single Euclidean one-sided Lipschitz constant, but a family of radial lower envelopes $r\mapsto\underline{\kappa}_{s}(r)$, one for each noise level $s$. The key point is that the same formula encodes both the near-field Euclidean obstruction and the far-field contraction reserve created by smoothing.
Definition 2 (Drift profile).
For a Borel vector field $u:\mathbb{R}^{d}\to\mathbb{R}^{d}$, define its radial drift profile by
$\kappa_{u}(r):=\inf_{\begin{subarray}{c}x\neq y\\ \|x-y\|=r\end{subarray}}-\frac{\langle u(x)-u(y),x-y\rangle}{\|x-y\|^{2}},\qquad r>0.$ (15)
Positive $\kappa_{u}(r)$ means contraction at separation $r$, while negative $\kappa_{u}(r)$ means expansion.
Gaussian smoothing transfers the weak-log-concavity parameters $(\alpha,M)$ of $p_{0}$ into the time-dependent quantities
$\alpha_{s}:=\frac{\alpha}{a^{2}(s)+\alpha\,\sigma^{2}(s)},\qquad M_{s}:=\frac{M\,a^{2}(s)}{\bigl(a^{2}(s)+\alpha\,\sigma^{2}(s)\bigr)^{2}},\qquad s\in[0,T].$ (16)
Here $\alpha_{s}$ quantifies the effective large-scale convexity of $p_{s}$, while $M_{s}$ measures the residual short-scale nonconvexity after smoothing.
Definition 3.
For $s\in(0,T]$, define
$\mathfrak{m}(s):=g^{2}(s)\bigl(\alpha_{s}-\ell(s)\bigr)-f(s),\qquad\mathfrak{b}(s):=f(s)+g^{2}(s)\bigl(M_{s}+\ell(s)-\alpha_{s}\bigr),$ (17)
and the radial lower envelope
$\underline{\kappa}_{s}(r):=g^{2}(s)\Big(\alpha_{s}-\frac{1}{r}f_{M_{s}}(r)-\ell(s)\Big)-f(s),\qquad r>0.$ (18)
For each fixed $s$, the same lower envelope can be rewritten as
$\underline{\kappa}_{s}(r)=\mathfrak{m}(s)-g^{2}(s)\frac{1}{r}f_{M_{s}}(r)=-\mathfrak{b}(s)+g^{2}(s)\Big(M_{s}-\frac{1}{r}f_{M_{s}}(r)\Big),\qquad r>0.$ (19)
This identity makes the geometry explicit: $\mathfrak{m}(s)$ is the far-field reserve, while the radius-dependent penalty $g^{2}(s)f_{M_{s}}(r)/r$ is largest near the origin and vanishes at infinity. Hence the reserve is converted into the Euclidean load $\mathfrak{b}(s)$ at short scales and reappears at large scales.
Lemma 1.
Under Assumptions 1 and 3, for $t\in[0,T]$ and s:=Tts:=T-t, the following hold:
  1. (i)
    Profile transfer. For every $r>0$,
    $\kappa_{\widehat{b}_{t}}(r)\geq\underline{\kappa}_{s}(r).$ (20)
  2. (ii)
    Endpoint structure. For each fixed $s\in(0,T]$, the map $r\mapsto\underline{\kappa}_{s}(r)$ is nondecreasing on $(0,\infty)$, with
    $\lim_{r\downarrow 0}\underline{\kappa}_{s}(r)=-\mathfrak{b}(s),\qquad\lim_{r\to\infty}\underline{\kappa}_{s}(r)=\mathfrak{m}(s).$ (21)
    If $\mathfrak{m}(s)>0$, then
    $R(s):=\inf\{r>0:\ \underline{\kappa}_{s}(r)\geq 0\}<\infty,$ (22)
    and $\underline{\kappa}_{s}(r)\geq 0$ for all $r\geq R(s)$.
  3. (iii)
    Companion Euclidean load. For every $x,y\in\mathbb{R}^{d}$,
    $\langle x-y,\widehat{b}_{t}(x)-\widehat{b}_{t}(y)\rangle\leq\mathfrak{b}(s)\|x-y\|^{2}.$ (23)
[…]
Figure 1: (a) Pointwise radial lower profile at a fixed noise level $s$. (b) Window-level quantities used by the theorem: $\underline{m}(s_{0})=\inf_{u\in[s_{0},T]}\mathfrak{m}(u)$ and $\Gamma(s_{0})=\int_{0}^{s_{0}}\mathfrak{b}(u)\,\mathop{}\!\mathrm{d}u$. Thus $s_{0}$ is admissible exactly when $\underline{m}(s_{0})>0$. The right panel is schematic and does not represent the evolution of $R(s)$.
The left panel is pointwise in $s$, whereas the right panel illustrates the uniform-in-time far-field positivity that underlies admissible switches.
Definition 4 (Admissible switches).
For $s\in[0,T]$ and $s_{0}\in(0,T]$, define
$\Gamma(s):=\int_{0}^{s}\mathfrak{b}(u)\,\mathop{}\!\mathrm{d}u,\qquad\underline{m}(s_{0}):=\inf_{u\in[s_{0},T]}\mathfrak{m}(u),$
whenever the integral defining $\Gamma(s)$ is finite. We call $s_{0}$ an admissible switch if $\underline{m}(s_{0})>0$, and write
$\mathcal{S}_{\mathrm{adm}}:=\{\,s_{0}\in(0,T]:\underline{m}(s_{0})>0\,\}.$
The infimum $\underline{m}(s_{0})$ enforces a positive far-field margin throughout the entire early window $s_{0},T]$. Thus admissibility does not mean that direct Euclidean closure is already available; it certifies only that a contractive radial regime is present before the return to $W_{2}$. For each admissible switch $s_{0}$, the appendices construct the associated switch metric $W_{\varphi_{s_{0}}}$, its damping rate $c(s_{0})$, and its affine-tail slope $\mathfrak{a}(s_{0})$.
4 A routing theorem from radial metric mismatch
Section 3 produced the theorem input from a single object: the radial lower profile $\underline{\kappa}_{s}$. Its far-field endpoint yields the admissible switch set $\mathcal{S}_{\mathrm{adm}}$, and its near-field endpoint yields the Euclidean load $\Gamma$. The theorem below turns this one-profile, two-endpoint structure into a one-switch routing bound.
For an admissible switch $s_{0}$, write
$\underline{g}(s_{0}):=\inf_{u\in[s_{0},T]}g(u),\qquad\overline{\mathfrak{b}}(s_{0}):=\sup_{u\in[s_{0},T]}\mathfrak{b}(u),\qquad\overline{G}(s_{0}):=\sup_{u\in[s_{0},T]}g^{2}(u)\sqrt{M_{u}}.$
For any increasing concave function $\varphi:[0,\infty)\to[0,\infty)$ with $\varphi(0)=0$, define the associated transport cost
$W_{\varphi}(\mu,\nu):=\inf_{\pi\in\Pi(\mu,\nu)}\int_{\mathbb{R}^{d}\times\mathbb{R}^{d}}\varphi(\|x-y\|)\,\pi(\mathop{}\!\mathrm{d}x,\mathop{}\!\mathrm{d}y).$
4.1 Main theorem
Throughout this subsection, work under Assumptions 1, 2, and 3. Let $0=t_{0}<\cdots<t_{N}=T$ be a numerical grid, and let $(Z_{t_{k}})_{k=0}^{N}$ be a Markovian discretization of the learned reverse SDE initialized by
$Z_{t_{0}}\sim\widehat{p}_{T},$
with numerical one-step kernels $\Psi_{k}$, i.e.
$\operatorname{Law}(Z_{t_{k+1}}\mid Z_{t_{k}}=x)=\Psi_{k}(x,\cdot),\qquad x\in\mathbb{R}^{d},\quad k=0,\dots,N-1.$
Write $P_{k}:=P_{t_{k},t_{k+1}}$ for the exact learned reverse one-step kernels. Assume that the ideal reverse SDE (12) admits a global strong solution $(\widetilde{X}_{t})_{0\leq t\leq T}$ on $[0,T]$ such that
$\operatorname{Law}(\widetilde{X}_{t})=p_{T-t},\qquad t\in[0,T].$
Assume moreover that the one-step defects satisfy
$W_{2}\!\bigl(P_{k}(x,\cdot),\Psi_{k}(x,\cdot)\bigr)\leq\xi_{k}(x),\qquad x\in\mathbb{R}^{d},\quad k=0,\dots,N-1,$ (24)
for Borel measurable functions $\xi_{k}:\mathbb{R}^{d}\to[0,\infty)$.
Theorem 1.
Fix $p>2$ and assume $\widehat{p}_{T}\in\mathcal{P}_{p}(\mathbb{R}^{d})$. Let $s_{0}\in\mathcal{S}_{\mathrm{adm}}$ be a grid-aligned admissible switch, set $t_{s}:=T-s_{0}=t_{K}$, and let $\varphi_{s_{0}}$, $c(s_{0})$, and $\mathfrak{a}(s_{0})$ denote the associated switch metric, early damping rate, and affine-tail slope. Assume that the learned reverse SDE is strongly well posed on $[0,T]$. Assume moreover that, for every $u,v]\subseteq[0,t_{s}]$ and every initial coupling $\pi\in\Pi(\mu,\nu)$ with $\mu,\nu\in\mathcal{P}_{1}(\mathbb{R}^{d})$, there exists a coalescing reflection coupling on $[u,v]$ with initial law $\pi$. Suppose further that
$\underline{g}(s_{0})>0,\qquad\overline{\mathfrak{b}}(s_{0})<\infty,\qquad\overline{G}(s_{0})<\infty,\qquad\Gamma(s_{0})<\infty.$
Let $\overline{\mathcal{M}}_{p}^{\mathrm{sw}}(s_{0})<\infty$ satisfy
$\overline{\mathcal{M}}_{p}^{\mathrm{sw}}(s_{0})\geq\mathop{\mathbb{E}}\nolimits\|Z_{t_{s}}\|^{p}+\mathop{\mathbb{E}}\nolimits_{X\sim p_{s_{0}}}\|X\|^{p}.$ (25)
Then
$W_{2}\!\bigl(\operatorname{Law}(Z_{t_{N}}),p_{0}\bigr)\leq\Delta_{\mathrm{late}}(s_{0})+e^{\Gamma(s_{0})}\,C_{p}^{\mathrm{sw}}(s_{0})\,\Bigl(\overline{\Delta}_{\varphi}^{\mathrm{sw}}(s_{0})\Bigr)^{\theta_{p}},$ (26)
where
$\theta_{p}:=\frac{p-2}{2(p-1)},\qquad C_{p}^{\mathrm{sw}}(s_{0}):=\sqrt{2(p-1)}(p-2)^{-\theta_{p}}\,\mathfrak{a}(s_{0})^{-\theta_{p}}\,\bigl(\overline{\mathcal{M}}_{p}^{\mathrm{sw}}(s_{0})\bigr)^{\frac{1}{2(p-1)}},$
and
$\overline{\Delta}_{\varphi}^{\mathrm{sw}}(s_{0}):=e^{-c(s_{0})t_{s}}\,W_{\varphi_{s_{0}}}(\widehat{p}_{T},p_{T})$ $+\sum_{k=0}^{K-1}e^{-c(s_{0})(t_{s}-t_{k+1})}\Bigl(\mathop{\mathbb{E}}\nolimits[\xi_{k}(Z_{t_{k}})^{2}]\Bigr)^{1/2}$
$+\int_{s_{0}}^{T}e^{-c(s_{0})(s-s_{0})}\,g^{2}(s)\,\varepsilon(s)\,\mathop{}\!\mathrm{d}s,$ (27)
$\Delta_{\mathrm{late}}(s_{0}):=\sum_{k=K}^{N-1}e^{\Gamma(T-t_{k+1})}\bigl(\mathop{\mathbb{E}}\nolimits\xi_{k}(Z_{t_{k}})^{2}\bigr)^{1/2}$ $+\int_{0}^{s_{0}}e^{\Gamma(s)}\,g^{2}(s)\,\varepsilon(s)\,\mathop{}\!\mathrm{d}s.$
Theorem 1 isolates a single geometric transition at $s_{0}$: before the interface, error is transported in the first geometry that actually contracts; after the interface, it is propagated in the Euclidean geometry in which the terminal discrepancy is measured. Thus $\overline{\Delta}_{\varphi}^{\mathrm{sw}}(s_{0})$ is a damped pre-switch budget, and $\theta_{p}=(p-2)/(2(p-1))$ is the one-time price of returning from the softer switch metric to $W_{2}$. Equivalently, the theorem induces a scalar switch-selection objective on the admissible grid: minimize the right-hand side of (26) over admissible grid-aligned $s_{0}$; see Corollary 2. Appendix D.2 shows in Proposition˜3 that this conversion price cannot be improved uniformly within the affine-tail concave class under the same switch-moment assumption.
The advantage is entirely an early-window phenomenon. Under weak log-concavity, the radial profile $\underline{\kappa}_{s}$ can become contractive in the far field before direct Euclidean closure is available. Reflection coupling sees this profile through the radial separation process, and the adapted metric $W_{\varphi_{s_{0}}}$ converts that far-field reserve into genuine transport contraction while absorbing the near-field obstruction. For fixed $s_{0}$, the routed and direct bounds share the same late-window term, so the comparison reduces exactly to the early window. The direct route weights an error injected at noise level $u\in[s_{0},T]$ by the residual Euclidean factor $e^{\int_{s_{0}}^{u}\mathfrak{b}(v)\,\mathop{}\!\mathrm{d}v}$, whereas the routed route first damps the same input by $e^{-c(s_{0})(u-s_{0})}$ and returns to $W_{2}$ only once. Hence the routed bound can be strictly sharper on mismatch windows. Appendix F.2 makes the comparison exact, and Appendix G.4 makes it concrete in a closed-form variance-preserving example.
5 Conclusion
The main message of this paper is that, under weak log-concavity, end-to-end error in reverse diffusion should not be propagated in a geometry chosen a priori. The central issue is not whether $W_{2}$ is the final error metric, but whether it is the right geometry in which to transport error at every stage of the reverse trajectory. The resulting gain is therefore not only geometric but quantitative: direct full-horizon Euclidean propagation can be genuinely conservative on mismatch windows.
More broadly, our result suggests that reverse-diffusion theory under weak regularity should be organized around geometry-adaptive error propagation rather than uniform metric propagation. The present paper isolates this principle in the simplest setting where it can be proved cleanly: a single radial interface and theorem-level bounds. A natural next step is to understand whether the same principle persists beyond one switch, beyond radial geometries, under weaker verifiable conditions, and in complexity guarantees tailored to concrete samplers.
Appendix roadmap.
The appendix follows the proof spine of the paper. Appendix˜A extracts the forward-time smoothing input and derives the quantities $\alpha_{s}$, $M_{s}$, $\mathfrak{m}(s)$, and $\mathfrak{b}(s)$. Appendix˜B transfers this profile to the learned reverse drift and characterizes admissible switches. Appendix˜C proves early-window contraction in the switch metric and packages the pre-switch initialization, discretization, and score-forcing terms. Appendix˜D proves the one-time return from $W_{\varphi_{s_{0}}}$ to $W_{2}$, including the sharpness of $\theta_{p}$. Appendix˜E propagates the converted discrepancy through the late Euclidean window, and Appendix˜F combines the pieces into the main theorem and the routed-versus-direct comparison. Appendix˜G records auxiliary comparison principles, conservative admissibility certificates, sufficient conditions for the early coupling input, and the explicit variance-preserving comparison result.
Appendix A Reverse-diffusion background and Gaussian smoothing
This section isolates the only forward-time regularization input used later. All subsequent quantities $\alpha_{s}$, $M_{s}$, $\mathfrak{m}(s)$, and $\mathfrak{b}(s)$ ultimately come from the representation
$p_{s}=(S_{a(s)}p_{0})*\gamma_{\sigma^{2}(s)}.$
For the theorem, it is not enough to know qualitatively that Gaussian smoothing improves geometry. We need the exact way in which smoothing acts on the weak profile
$\kappa_{V_{0}}(r)\geq\alpha-\frac{1}{r}f_{M}(r).$
There are three operations behind the formulas proved below: deterministic scaling by $a(s)$, which weakens both curvature and defect by the factor $a(s)^{-2}$; Gaussian convolution, which preserves the weak semiconvex defect once the available quadratic part has been peeled off; and a quadratic tilt, which returns part of that peeled-off curvature in explicit form. The outcome is the pair $(\alpha_{s},M_{s})$: $\alpha_{s}$ records the large-scale convexity that survives smoothing, while $M_{s}$ records the residual short-scale defect that smoothing has not yet erased. Nothing in this section depends on the later metric switch. Its outputs are the explicit profile bound for $V_{s}:=-\log p_{s}$ and the moment bounds used at switch time.
Throughout, we use the drift profile from Definition 2. When $h\in C^{1}(\mathbb{R}^{d})$ is a potential, we write
$\kappa_{h}(r):=\kappa_{-\nabla h}(r)=\inf_{\begin{subarray}{c}x\neq y\\ \|x-y\|=r\end{subarray}}\frac{\langle\nabla h(x)-\nabla h(y),x-y\rangle}{\|x-y\|^{2}},\qquad r>0.$
Thus $\kappa_{h}$ is simply the profile of the gradient drift $\nabla h$.
A.1 Forward smoothing and basic regularity
Lemma 2.
For every $s\in(0,T]$, one has $\sigma(s)>0$. Consequently, $p_{s}=\operatorname{Law}(X_{s})$ admits a strictly positive $C^{\infty}$ density, and
$V_{s}:=-\log p_{s}\in C^{\infty}(\mathbb{R}^{d}).$
In particular, $\kappa_{V_{s}}$ is well defined for every $s>0$.
Proof.
Since $g(u)>0$ for $u>0$, the representation
$\sigma^{2}(s)=\int_{0}^{s}\exp\!\Big(-2\int_{u}^{s}f(v)\,\mathop{}\!\mathrm{d}v\Big)\,g^{2}(u)\,\mathop{}\!\mathrm{d}u$
has strictly positive integrand, hence $\sigma(s)>0$ for $s>0$. Moreover,
$X_{s}\stackrel{{\scriptstyle d}}{{=}}a(s)X_{0}+\sigma(s)\xi,\qquad\xi\sim\mathcal{N}(0,I_{d}),$
so $p_{s}=(S_{a(s)}p_{0})*\mathcal{N}(0,\sigma^{2}(s)I_{d})$. Thus $p_{s}$ is the convolution of a probability measure with a nondegenerate Gaussian density, hence it is strictly positive and $C^{\infty}$. Therefore $V_{s}=-\log p_{s}\in C^{\infty}(\mathbb{R}^{d})$. ∎
A.2 Profile calculus under OU smoothing
The next bounds package the smoothing mechanism above into the form used later. The scaling lemma isolates what the deterministic factor $a(s)$ does to the profile. The weak-semiconvexity invariance of Gaussian convolution from Conforti [2024, Theorem 2.1] is then applied not to $U$ itself but to $h=U-\frac{\alpha^{\prime}}{2}\|\cdot\|^{2}$, because the defect class $\mathcal{F}_{M}$ is stable only after the explicit quadratic curvature has been separated off. The factorization lemma explains why convolving $e^{-U}$ with a Gaussian becomes a heat flow of $e^{-h}$ evaluated at a rescaled point, multiplied by a new quadratic tilt. That is the precise origin of the denominator $1+\alpha^{\prime}\sigma^{2}$ in the final profile formulas; see also Gentiloni-Silveri and Ocello [2025, Appendix B] for the same input in the diffusion-model setting.
Lemma 3.
For all $M>0$ and $r\geq 0$,
$0\leq f_{M}(r)\leq\min\{Mr,\ 2\sqrt{M}\}.$
Consequently, for every $r>0$,
$\frac{1}{r}f_{M}(r)\leq\min\left\{M,\ \frac{2\sqrt{M}}{r}\right\}.$
Proof.
The claim is trivial if $M=0$ or $r=0$. Otherwise, with $y=\frac{1}{2}\sqrt{M}\,r$, one has $f_{M}(r)=2\sqrt{M}\,\tanh(y)$, and $\tanh(y)\leq\min\{y,1\}$ gives
$f_{M}(r)\leq\min\{2\sqrt{M}\,y,\ 2\sqrt{M}\}=\min\{Mr,\ 2\sqrt{M}\}.$
Dividing by $r>0$ yields the second bound. ∎
For $u>0$, let $\gamma_{u}$ denote the density of $\mathcal{N}(0,uI_{d})$, and define
$(P_{u}\phi)(x):=\int_{\mathbb{R}^{d}}\phi(y)\,\gamma_{u}(x-y)\,\mathop{}\!\mathrm{d}y.$
For $M>0$, set
$\mathcal{F}_{M}:=\Bigl\{h\in C^{1}(\mathbb{R}^{d}):\ \kappa_{h}(r)\geq-\frac{1}{r}f_{M}(r)\ \ \forall r>0\Bigr\}.$
Lemma 4.
Let $p_{0}\propto e^{-V_{0}}$ with $V_{0}\in C^{1}(\mathbb{R}^{d})$, and for $a>0$ let
$q:=S_{a}p_{0},\qquad U_{a}:=-\log q.$
Then for every $r>0$,
$\kappa_{U_{a}}(r)=\frac{1}{a^{2}}\,\kappa_{V_{0}}(r/a).$
In particular, if
$\kappa_{V_{0}}(r)\geq\alpha-\frac{1}{r}f_{M}(r),\qquad r>0,$
then
$\kappa_{U_{a}}(r)\geq\frac{\alpha}{a^{2}}-\frac{1}{r}f_{M/a^{2}}(r),\qquad r>0.$
Proof.
Since $q(x)=a^{-d}p_{0}(x/a)$, one has $U_{a}(x)=V_{0}(x/a)+\mathrm{const}$ and $\nabla U_{a}(x)=a^{-1}\nabla V_{0}(x/a)$. Thus, for $u=x/a$, $v=y/a$,
$\frac{\langle\nabla U_{a}(x)-\nabla U_{a}(y),x-y\rangle}{\|x-y\|^{2}}=\frac{1}{a^{2}}\frac{\langle\nabla V_{0}(u)-\nabla V_{0}(v),u-v\rangle}{\|u-v\|^{2}}.$
Taking the infimum over $\|x-y\|=r$, equivalently $\|u-v\|=r/a$, gives the first identity. The second follows by applying the assumed lower bound at radius $r/a$ and using $a^{-1}f_{M}(r/a)=f_{M/a^{2}}(r)$. ∎
Lemma 5 (Heat semigroup invariance of $\mathcal{F}_{M}$).
Let $M>0$ and $h\in\mathcal{F}_{M}$. For $u>0$, define
$H_{u}(x):=-\log\big(P_{u}(e^{-h})\big)(x).$
Then $H_{u}\in\mathcal{F}_{M}$.
Proof.
This is exactly Conforti [2024, Theorem 2.1]. ∎
Lemma 6.
Let $U:\mathbb{R}^{d}\to\mathbb{R}$, $\sigma>0$, and $\alpha^{\prime}\geq 0$. Define
$h(y):=U(y)-\frac{\alpha^{\prime}}{2}\|y\|^{2},\qquad A:=1+\alpha^{\prime}\sigma^{2},\qquad\tau:=\frac{\sigma^{2}}{A}.$
Let
$p(x):=\int_{\mathbb{R}^{d}}e^{-U(y)}\,\gamma_{\sigma^{2}}(x-y)\,\mathop{}\!\mathrm{d}y.$
Then there exists $C=C(\alpha^{\prime},\sigma,d)$ such that for all $x\in\mathbb{R}^{d}$,
$p(x)=C\exp\!\Big(-\frac{\alpha^{\prime}}{2A}\|x\|^{2}\Big)\Big(P_{\tau}(e^{-h})\Big)\!\Big(\frac{x}{A}\Big).$
Proof.
Since $e^{-U(y)}=e^{-h(y)}e^{-\alpha^{\prime}\|y\|^{2}/2}$, one gets
$p(x)=(2\pi\sigma^{2})^{-d/2}\int_{\mathbb{R}^{d}}e^{-h(y)}\exp\!\Big(-\frac{\alpha^{\prime}}{2}\|y\|^{2}-\frac{\|x-y\|^{2}}{2\sigma^{2}}\Big)\,\mathop{}\!\mathrm{d}y.$
Completing the square gives
$\frac{\alpha^{\prime}}{2}\|y\|^{2}+\frac{\|x-y\|^{2}}{2\sigma^{2}}=\frac{\alpha^{\prime}}{2A}\|x\|^{2}+\frac{1}{2\tau}\Big\|y-\frac{x}{A}\Big\|^{2}.$
Recognizing the Gaussian density $\gamma_{\tau}$ yields
$p(x)=A^{-d/2}\exp\!\Big(-\frac{\alpha^{\prime}}{2A}\|x\|^{2}\Big)\Big(P_{\tau}(e^{-h})\Big)\!\Big(\frac{x}{A}\Big),$
so the claim holds with $C=A^{-d/2}$. ∎
Lemma 7.
Let $\alpha^{\prime}\geq 0$ and $M^{\prime}\geq 0$. Assume $q\propto e^{-U}$ with $U\in C^{1}(\mathbb{R}^{d})$ satisfies
$\kappa_{U}(r)\geq\alpha^{\prime}-\frac{1}{r}f_{M^{\prime}}(r),\qquad r>0.$
Let $p=q*\gamma_{\sigma^{2}}$ and $V:=-\log p$. Then for all $r>0$,
$\kappa_{V}(r)\geq\frac{\alpha^{\prime}}{1+\alpha^{\prime}\sigma^{2}}-\frac{1}{r}f_{\,M^{\prime}/(1+\alpha^{\prime}\sigma^{2})^{2}}(r).$
Proof.
The proof separates the explicit quadratic curvature from the residual defect. The residual part belongs to a class invariant under Gaussian convolution, while the quadratic part can be tracked exactly by completing the square. Set $A:=1+\alpha^{\prime}\sigma^{2}$, $\tau:=\sigma^{2}/A$, and $h:=U-\frac{\alpha^{\prime}}{2}\|\cdot\|^{2}$. Then $\kappa_{h}(r)=\kappa_{U}(r)-\alpha^{\prime}$, so $h\in\mathcal{F}_{M^{\prime}}$. By Lemma 6,
$V(x)=\frac{\alpha^{\prime}}{2A}\|x\|^{2}+H(x/A)+\mathrm{const},\qquad H:=-\log\big(P_{\tau}(e^{-h})\big).$
Lemma 5 gives $H\in\mathcal{F}_{M^{\prime}}$. Using the quadratic contribution and the same scaling computation as in Lemma 4,
$\kappa_{V}(r)\geq\frac{\alpha^{\prime}}{A}+\frac{1}{A^{2}}\kappa_{H}(r/A)\geq\frac{\alpha^{\prime}}{A}-\frac{1}{Ar}f_{M^{\prime}}(r/A)=\frac{\alpha^{\prime}}{A}-\frac{1}{r}f_{M^{\prime}/A^{2}}(r).$
Corollary 1 (OU smoothing of the weak profile).
Assume $p_{0}\propto e^{-V_{0}}$ satisfies Assumption 1 with parameters $(\alpha,M)$. For $s\in(0,T]$, let
$p_{s}=(S_{a(s)}p_{0})*\gamma_{\sigma^{2}(s)},\qquad V_{s}:=-\log p_{s}.$
Then for every $r>0$,
$\kappa_{V_{s}}(r)\geq\alpha_{s}-\frac{1}{r}f_{M_{s}}(r),$
where
$\alpha_{s}=\frac{\alpha}{a^{2}(s)+\alpha\,\sigma^{2}(s)},\qquad M_{s}=\frac{M\,a^{2}(s)}{(a^{2}(s)+\alpha\,\sigma^{2}(s))^{2}}.$
Proof.
Let $U_{s}$ denote the potential of $S_{a(s)}p_{0}$. Lemma 4 gives
$\kappa_{U_{s}}(r)\geq\frac{\alpha}{a^{2}(s)}-\frac{1}{r}f_{M/a^{2}(s)}(r).$
Applying Lemma 7 with $\alpha^{\prime}=\alpha/a^{2}(s)$, $M^{\prime}=M/a^{2}(s)$, and $\sigma=\sigma(s)$ yields
$\kappa_{V_{s}}(r)\geq\frac{\alpha/a^{2}(s)}{1+(\alpha/a^{2}(s))\sigma^{2}(s)}-\frac{1}{r}f_{\frac{M/a^{2}(s)}{(1+(\alpha/a^{2}(s))\sigma^{2}(s))^{2}}}(r),$
which simplifies to the stated formulas for $\alpha_{s}$ and $M_{s}$. ∎
Remark 2 (Why $\alpha_{s}$ and $M_{s}$ share the same denominator).
The common denominator $a^{2}(s)+\alpha\,\sigma^{2}(s)$ is the algebraic signature of one smoothing mechanism acting on two different parts of the profile. The factor $a^{2}(s)$ comes from transporting the initial geometry through the linear part of the forward dynamics, while the term $\alpha\,\sigma^{2}(s)$ is the quadratic curvature that can be fed into the Gaussian-convolution step. The numerators then separate the two roles. The large-scale convex part keeps the factor $\alpha$, producing $\alpha_{s}$, whereas the short-scale defect keeps the factor $Ma^{2}(s)$, producing $M_{s}$. Thus smoothing improves the available convexity and the residual defect through one common denominator, but only the defect remembers how much unsmoothed nonconvexity was still present before convolution.
A.3 Moment consequences
We also need one integrability fact later at switch time. The point is that the defect term in Assumption 1 is uniformly bounded: by Lemma 3, $r^{-1}f_{M}(r)$ can at worst create a linear loss along rays, while the positive $\alpha r$ term integrates to a quadratic confinement. Thus weak log-concavity with $\alpha>0$ already forces Gaussian-type tails.
Lemma 8.
Under Assumption 1 with parameters $\alpha,M>0$, there exists a constant $C_{0}<\infty$ such that
$V_{0}(x)\geq\frac{\alpha}{4}\|x\|^{2}-C_{0},\qquad\forall x\in\mathbb{R}^{d}.$
Consequently,
$p_{0}\in\mathcal{P}_{p}(\mathbb{R}^{d})\qquad\text{for every }p\geq 1.$
Moreover, all forward marginals $p_{s}=\operatorname{Law}(X_{s})$ also belong to $\mathcal{P}_{p}(\mathbb{R}^{d})$ for every $p\geq 1$.
Proof.
Fix $x=ru$ with $r=\|x\|>0$ and $\|u\|=1$. Applying Assumption 1 with $y=0$ gives
$\langle\nabla V_{0}(ru)-\nabla V_{0}(0),u\rangle\geq\alpha r-f_{M}(r),$
hence
$\langle\nabla V_{0}(ru),u\rangle\geq\alpha r-f_{M}(r)-\|\nabla V_{0}(0)\|.$
Integrating along the ray $s\mapsto su$,
$V_{0}(ru)-V_{0}(0)=\int_{0}^{r}\langle\nabla V_{0}(su),u\rangle\,\mathop{}\!\mathrm{d}s\geq\frac{\alpha}{2}r^{2}-\int_{0}^{r}f_{M}(s)\,\mathop{}\!\mathrm{d}s-\|\nabla V_{0}(0)\|\,r.$
By Lemma 3, $f_{M}(s)\leq 2\sqrt{M}$, so
$V_{0}(x)\geq V_{0}(0)+\frac{\alpha}{2}\|x\|^{2}-\bigl(2\sqrt{M}+\|\nabla V_{0}(0)\|\bigr)\|x\|.$
Absorbing the linear term into the quadratic term yields $V_{0}(x)\geq\frac{\alpha}{4}\|x\|^{2}-C_{0}$ for some $C_{0}<\infty$. Thus $e^{-V_{0}(x)}\lesssim e^{-\alpha\|x\|^{2}/4}$, so $p_{0}$ has finite moments of every order. The same then holds for $p_{s}$ because
$X_{s}\stackrel{{\scriptstyle d}}{{=}}a(s)X_{0}+\sigma(s)\xi,\qquad\xi\sim\mathcal{N}(0,I_{d}),$
and Gaussian moments are finite. ∎
Appendix B Profile transfer, Euclidean closure, and admissible switches
This section translates the smoothing output into geometry of the learned reverse drift. For each noise level $s$, the learned reverse drift inherits a radial lower envelope
$\underline{\kappa}_{s}(r)=\mathfrak{m}(s)-g^{2}(s)\frac{1}{r}f_{M_{s}}(r),\qquad r>0,$
where s=Tts=T-t. The key point is that this is one and the same monotone-in-$r$ curve. Its left endpoint is the Euclidean load $\mathfrak{b}(s)$, which is the quantity seen by synchronous coupling. Its right endpoint is the far-field reserve $\mathfrak{m}(s)$, which is the quantity later converted into contraction by reflection coupling. Thus the later coupling split is not imposed from outside: it is already encoded in the shape of $\underline{\kappa}_{s}$. No stochastic argument is needed yet; the whole section is deterministic radial bookkeeping.
B.1 Endpoint structure of the lower radial profile
Proof of Lemma 1.
Fix $t\in[0,T]$ and write s:=Tts:=T-t.
Item (i): profile transfer. For $x\neq y$, set $r:=\|x-y\|$. Since
$\widehat{b}_{t}(x)=f(s)x+g^{2}(s)\nabla\log p_{s}(x)+g^{2}(s)e_{s}(x),$
one has
$\frac{\langle\widehat{b}_{t}(x)-\widehat{b}_{t}(y),x-y\rangle}{\|x-y\|^{2}}=-f(s)-g^{2}(s)\frac{\langle\nabla\log p_{s}(x)-\nabla\log p_{s}(y),x-y\rangle}{\|x-y\|^{2}}-g^{2}(s)\frac{\langle e_{s}(x)-e_{s}(y),x-y\rangle}{\|x-y\|^{2}}.$
Because $\nabla\log p_{s}=-\nabla V_{s}$, the middle term is $\kappa_{V_{s}}(r)$. By Corollary 1,
$\kappa_{V_{s}}(r)\geq\alpha_{s}-\frac{1}{r}f_{M_{s}}(r),$
and Assumption 3 gives
$\frac{\langle e_{s}(x)-e_{s}(y),x-y\rangle}{\|x-y\|^{2}}\leq\ell(s).$
Combining the two bounds yields
$\frac{\langle\widehat{b}_{t}(x)-\widehat{b}_{t}(y),x-y\rangle}{\|x-y\|^{2}}\geq g^{2}(s)\Big(\alpha_{s}-\frac{1}{r}f_{M_{s}}(r)-\ell(s)\Big)-f(s)=\underline{\kappa}_{s}(r).$
Taking the infimum over all pairs with $\|x-y\|=r$ proves (20).
Item (ii): endpoint structure of the same profile. Define
$q_{M}(r):=\frac{1}{r}f_{M}(r),\qquad r>0.$
Then
$\underline{\kappa}_{s}(r)=\mathfrak{m}(s)-g^{2}(s)\,q_{M_{s}}(r).$
If $M_{s}=0$, then $f_{M_{s}}\equiv 0$, so $\underline{\kappa}_{s}(r)\equiv\mathfrak{m}(s)$, and all claims are immediate. Assume now that $M_{s}>0$. With $z=\frac{1}{2}\sqrt{M_{s}}\,r$ and $\psi(z):=\tanh(z)/z$, one has
$q_{M_{s}}(r)=M_{s}\,\psi\!\left(\frac{1}{2}\sqrt{M_{s}}\,r\right).$
Moreover,
$\psi^{\prime}(z)=\frac{z\,\operatorname{sech}^{2}(z)-\tanh(z)}{z^{2}}=-\frac{\tanh(z)-z\,\operatorname{sech}^{2}(z)}{z^{2}}.$
Set $h(z):=\tanh(z)-z\,\operatorname{sech}^{2}(z)$. Then $h(0)=0$ and
$h^{\prime}(z)=2z\,\operatorname{sech}^{2}(z)\tanh(z)\geq 0\qquad(z\geq 0),$
so $h(z)\geq 0$ for $z\geq 0$. Hence $\psi^{\prime}(z)\leq 0$, and therefore $r\mapsto q_{M_{s}}(r)$ is nonincreasing on $(0,\infty)$. Consequently $r\mapsto\underline{\kappa}_{s}(r)$ is nondecreasing on $(0,\infty)$.
The endpoint limits now follow from
$\lim_{z\downarrow 0}\frac{\tanh(z)}{z}=1,\qquad\lim_{z\to\infty}\frac{\tanh(z)}{z}=0.$
Indeed,
$\lim_{r\downarrow 0}q_{M_{s}}(r)=M_{s},\qquad\lim_{r\to\infty}q_{M_{s}}(r)=0,$
hence
$\lim_{r\downarrow 0}\underline{\kappa}_{s}(r)=\mathfrak{m}(s)-g^{2}(s)M_{s}=-\mathfrak{b}(s),$
and
$\lim_{r\to\infty}\underline{\kappa}_{s}(r)=\mathfrak{m}(s).$
The identity is just the rearrangement
$\mathfrak{m}(s)+\mathfrak{b}(s)=g^{2}(s)M_{s}.$
If $\mathfrak{m}(s)>0$, the second limit implies that there exists $r_{\ast}>0$ such that $\underline{\kappa}_{s}(r_{\ast})>0$. By monotonicity, $\underline{\kappa}_{s}(r)\geq 0$ for all $r\geq r_{\ast}$. Therefore the set in (22) is nonempty, so $R(s)<\infty$, and $\underline{\kappa}_{s}(r)\geq 0$ for all $r\geq R(s)$.
Item (iii): companion Euclidean load. By item (ii),
$\underline{\kappa}_{s}(r)\geq\lim_{u\downarrow 0}\underline{\kappa}_{s}(u)=-\mathfrak{b}(s),\qquad r>0.$
Together with item (i), this yields
$\kappa_{\widehat{b}_{t}}(r)\geq-\mathfrak{b}(s),\qquad r>0.$
Hence, for $x\neq y$,
$\langle x-y,\widehat{b}_{t}(x)-\widehat{b}_{t}(y)\rangle\leq-\kappa_{\widehat{b}_{t}}(\|x-y\|)\|x-y\|^{2}\leq\mathfrak{b}(s)\|x-y\|^{2}.$
The case $x=y$ is trivial, proving (23). ∎
Remark 3 (Geometry behind the later coupling split).
Fix $s\in(0,T]$. The curve $r\mapsto\underline{\kappa}_{s}(r)$ rises from the near-field value $\mathfrak{b}(s)$ to the far-field value $\mathfrak{m}(s)$. So Euclidean geometry and radial geometry are reading different parts of the same object. A synchronous-coupling estimate for $\|Y_{t}-\bar{Y}_{t}\|^{2}$ collapses the whole curve to its left endpoint, because it must control pairwise separation uniformly over all radii. By contrast, reflection coupling can act directly on the radial process and therefore exploit the fact that the same profile may already be positive in the tail even when its near-field endpoint is still negative. This is exactly the metric-mismatch regime: far-field contraction has appeared, but Euclidean closure has not. The switch metric introduced later is tailored to that shape—it suppresses the near-field defect while keeping sensitivity to the contractive tail.
B.2 Characterization of admissible switch times
Definition 4 asks that the far-field endpoint stay positive on the whole remaining noise interval. By Remark 3, this does not mean that Euclidean geometry has already been restored. It only means that every instantaneous profile on $s_{0},T]$ has a positive tail, which is the minimum robust input needed later to build one time-uniform switch metric and run a single reflection-coupling argument on the whole early window.
Proposition 1 (Structure of the admissible switch set).
Assume that $\mathfrak{m}$ is continuous on $(0,T]$. Then the map
$s_{0}\longmapsto\underline{m}(s_{0}):=\inf_{u\in[s_{0},T]}\mathfrak{m}(u)$
is nondecreasing and continuous on $(0,T]$. Consequently,
$\mathcal{S}_{\mathrm{adm}}=\{\,s_{0}\in(0,T]:\underline{m}(s_{0})>0\,\}$
is an upper interval of $(0,T]$. If, in addition, $\mathfrak{m}(T)>0$, then a nontrivial admissible switch exists. Moreover, if $s_{\min}:=\inf\mathcal{S}_{\mathrm{adm}}\in(0,T)$, then
$\mathfrak{m}(s_{\min})=0.$
Proof.
If $0<s_{1}\leq s_{2}\leq T$, then $s_{2},T]\subseteq[s_{1},T]$, hence
$\underline{m}(s_{1})=\inf_{u\in[s_{1},T]}\mathfrak{m}(u)\leq\inf_{u\in[s_{2},T]}\mathfrak{m}(u)=\underline{m}(s_{2}),$
so $s_{0}\mapsto\underline{m}(s_{0})$ is nondecreasing.
To prove continuity, define
$F(s_{0},\theta):=\mathfrak{m}\bigl((1-\theta)s_{0}+\theta T\bigr),\qquad(s_{0},\theta)\in(0,T]\times[0,1].$
Since $\mathfrak{m}$ is continuous, so is $F$, and
$\underline{m}(s_{0})=\min_{\theta\in[0,1]}F(s_{0},\theta).$
The minimum of a continuous function over the compact set $[0,1]$ depends continuously on $s_{0}$, hence $\underline{m}$ is continuous on $(0,T]$.
Because $\mathcal{S}_{\mathrm{adm}}=\underline{m}^{-1}((0,\infty))$ and $\underline{m}$ is nondecreasing, $\mathcal{S}_{\mathrm{adm}}$ is an upper interval. If $\mathfrak{m}(T)>0$, continuity gives $\eta>0$ such that $\mathfrak{m}(u)>0$ on $T-\eta,T]$, hence $\underline{m}(T-\eta)>0$, so a nontrivial admissible switch exists. Finally, let $s_{\min}:=\inf\mathcal{S}_{\mathrm{adm}}\in(0,T)$. Since $\mathcal{S}_{\mathrm{adm}}$ is an upper interval, every $u>s_{\min}$ belongs to $\mathcal{S}_{\mathrm{adm}}$, hence
$\underline{m}(u)>0,\qquad u\in(s_{\min},T].$
Therefore, for every $u>s_{\min}$ and every $r\in[u,T]$,
$\mathfrak{m}(r)\geq\underline{m}(u)>0.$
As $u>s_{\min}$ is arbitrary, this implies
$\mathfrak{m}(r)>0,\qquad r\in(s_{\min},T].$
Suppose, for contradiction, that $\mathfrak{m}(s_{\min})>0$. Then continuity of $\mathfrak{m}$ implies that $\mathfrak{m}$ is strictly positive on the compact interval $s_{\min},T]$, so
$\underline{m}(s_{\min})=\min_{u\in[s_{\min},T]}\mathfrak{m}(u)>0.$
By continuity of $\underline{m}$, there exists $\eta>0$ such that
$\underline{m}(s_{\min}-\eta)>0,$
contradicting the definition of $s_{\min}$. Thus $\mathfrak{m}(s_{\min})=0$. ∎
Remark 4 (Why admissibility is the right threshold for the next section).
Admissibility is intentionally weaker than Euclidean stability. It does not require $\mathfrak{b}(u)\leq 0$ on $s_{0},T]$, so one cannot generally replace the early radial route by a direct $W_{2}$ argument on that whole window. What it guarantees instead is that the positive tail survives uniformly after taking the infimum over $u\in[s_{0},T]$. The next section turns exactly that surviving tail positivity into a two-zone lower envelope with a bounded near-field loss and a strictly positive far-field reserve.
Remark 5 (The radial portrait in the main text is schematic).
Figure 1 should be read as a pointwise visualization of the lower profile $(s,r)\mapsto\underline{\kappa}_{s}(r)$. The threshold radius $R(s)$ is only an instantaneous zero-crossing scale. Later we will replace it by the switch-dependent radius $R_{\mathrm{sw}}(s_{0})$, which is a uniform envelope radius on the whole early window rather than a pointwise radius at a single $s$. No claim is made here that $R(s)$ moves monotonically inward with the noise level.
Appendix C Early-phase damping in the switch metric
Fix a grid-aligned admissible switch $s_{0}\in\mathcal{S}_{\mathrm{adm}}$, and write
$t_{s}:=T-s_{0}=t_{K}.$
This section proves the early-window contraction used later. The argument has three steps. First, we replace the family
$\bigl(r\mapsto\underline{\kappa}_{s}(r)\bigr)_{s\in[s_{0},T]}$
by a single switch-dependent lower envelope that is uniform on $0,t_{s}]$. Second, we use reflection coupling for two copies of the learned reverse SDE to obtain semigroup contraction in $W_{\varphi_{s_{0}}}$. Third, we insert the three pre-switch error channels—initialization, discretization, and score forcing—at their entry times and propagate them to $t_{s}$ under that contraction.
Two couplings are used, for different reasons. Reflection coupling is used only for the learned-semigroup contraction. By contrast, when we compare the ideal and learned reverse dynamics on a short interval, we use synchronous coupling so that the noise cancels and only the drift mismatch $\delta_{t}$ remains. Throughout this section, $P_{u,v}$ denotes the exact learned reverse semigroup on $[u,v]$, $P_{k}:=P_{t_{k},t_{k+1}}$ the exact learned one-step kernels, and $\Psi_{k}$ the numerical one-step kernels; we write
$\delta_{t}(x):=\widehat{b}_{t}(x)-b_{t}(x)=g^{2}(T-t)e_{T-t}(x).$ (28)
From this point on, two theorem-level inputs are used repeatedly. Whenever Lemma 13 is invoked, we use strong well posedness of the learned reverse SDE on $0,t_{s}]$, existence of coalescing reflection couplings on every subinterval, and the finiteness conditions
$\underline{g}(s_{0})>0,\qquad\overline{\mathfrak{b}}(s_{0})<\infty,\qquad\overline{G}(s_{0})<\infty.$
Whenever Lemma 14 is invoked, we also use the existence of a global ideal reverse solution
$(\widetilde{X}_{t})_{0\leq t\leq T}$
to (12) with
$\operatorname{Law}(\widetilde{X}_{t})=p_{T-t},\qquad 0\leq t\leq T.$
C.1 A uniform early-window profile from admissibility
The first deliberate loss of information occurs here. We replace the full family $\bigl(r\mapsto\underline{\kappa}_{s}(r)\bigr)_{s\in[s_{0},T]}$ by a switch-dependent two-zone lower envelope: a bounded near-field load and a strictly positive far-field reserve. This loses the detailed radius dependence of the exact profile, but gains the uniformity in $t$ needed by the later coupling argument.
Define
$\underline{g}(s_{0}):=\inf_{u\in[s_{0},T]}g(u),\qquad\overline{\mathfrak{b}}(s_{0}):=\sup_{u\in[s_{0},T]}\mathfrak{b}(u),\qquad\overline{G}(s_{0}):=\sup_{u\in[s_{0},T]}g^{2}(u)\sqrt{M_{u}},$
$R_{\mathrm{sw}}(s_{0}):=\frac{4\,\overline{G}(s_{0})}{\underline{m}(s_{0})},\qquad m_{\mathrm{sw}}(s_{0}):=\frac{\underline{m}(s_{0})}{2}.$
Here $R_{\mathrm{sw}}(s_{0})$ is a switch-dependent uniform envelope radius on the whole early window; it is not the instantaneous zero-crossing radius $R(s)$ from (22). Define the switch-level lower profile
$\underline{\kappa}_{s_{0}}(r):=\begin{cases}-\overline{\mathfrak{b}}(s_{0}),&0<r\leq R_{\mathrm{sw}}(s_{0}),\\ m_{\mathrm{sw}}(s_{0}),&r>R_{\mathrm{sw}}(s_{0}).\end{cases}$
Lemma 9.
Fix $s_{0}\in\mathcal{S}_{\mathrm{adm}}$ and assume $\overline{\mathfrak{b}}(s_{0})<\infty$ and $\overline{G}(s_{0})<\infty$. Then for every $t\in[0,t_{s}]$ and every $r>0$,
$\kappa_{\widehat{b}_{t}}(r)\geq\underline{\kappa}_{s_{0}}(r),$ (29)
and
$g(T-t)\geq\underline{g}(s_{0}).$ (30)
Proof.
Fix $t\in[0,t_{s}]$ and set $s:=T-t\in[s_{0},T]$. By Lemma 1(i),
$\kappa_{\widehat{b}_{t}}(r)\geq\underline{\kappa}_{s}(r)=\mathfrak{m}(s)-g^{2}(s)\frac{1}{r}f_{M_{s}}(r),\qquad r>0.$
If $0<r\leq R_{\mathrm{sw}}(s_{0})$, then Lemma 3 gives $r^{-1}f_{M_{s}}(r)\leq M_{s}$, hence
$\kappa_{\widehat{b}_{t}}(r)\geq g^{2}(s)\bigl(\alpha_{s}-\ell(s)-M_{s}\bigr)-f(s)=-\mathfrak{b}(s)\geq-\overline{\mathfrak{b}}(s_{0}).$
If $r>R_{\mathrm{sw}}(s_{0})$, then Lemma 3 gives $f_{M_{s}}(r)\leq 2\sqrt{M_{s}}$, and therefore
$\kappa_{\widehat{b}_{t}}(r)\geq\mathfrak{m}(s)-\frac{2g^{2}(s)\sqrt{M_{s}}}{r}\geq\underline{m}(s_{0})-\frac{2\overline{G}(s_{0})}{r}.$
By the definition of $R_{\mathrm{sw}}(s_{0})$, the last term is bounded below by
$\underline{m}(s_{0})-\frac{2\overline{G}(s_{0})}{R_{\mathrm{sw}}(s_{0})}=\frac{\underline{m}(s_{0})}{2}=m_{\mathrm{sw}}(s_{0}).$
This proves (29). The bound (30) is immediate from the definition of $\underline{g}(s_{0})$. ∎
C.2 Reflection coupling and radial reduction
With the uniform profile in hand, the problem becomes one-dimensional. For two copies of the learned reverse SDE, reflection coupling is the natural choice because it puts the stochastic forcing entirely in the radial direction of the separation vector. The distance process then sees the drift only through the scalar profile $\kappa_{\widehat{b}_{t}}(r_{t})$, exactly the object extracted in the previous subsection.
Definition 5 (Coalescing reflection coupling).
Fix $0\leq t_{1}<t_{2}\leq t_{s}$. Let $Y$ and $\bar{Y}$ be solutions of the learned reverse SDE (13) on $t_{1},t_{2}]$, defined on a common filtered probability space. Set
$D_{t}:=Y_{t}-\bar{Y}_{t},\qquad r_{t}:=\|D_{t}\|,\qquad\tau_{c}:=\inf\{t\geq t_{1}:\ r_{t}=0\}.$
Fix a deterministic unit vector $e\in\mathbb{R}^{d}$, and define
$n_{t}:=\frac{D_{t}}{r_{t}}\mathbf{1}_{\{r_{t}>0\}}+e\,\mathbf{1}_{\{r_{t}=0\}},\qquad H_{t}:=(I_{d}-2n_{t}n_{t}^{\top})\mathbf{1}_{\{t<\tau_{c}\}}+I_{d}\,\mathbf{1}_{\{t\geq\tau_{c}\}}.$
If $Y$ is driven by a Brownian motion $\overline{B}$, define
$\overline{B}_{t}^{\,\mathrm{ref}}:=\int_{t_{1}}^{t}H_{u}\,\mathop{}\!\mathrm{d}\overline{B}_{u}.$
The pair $(Y,\bar{Y})$ is said to be in coalescing reflection coupling if $Y$ is driven by $\overline{B}$ and $\bar{Y}$ is driven by $\overline{B}^{\,\mathrm{ref}}$. Since $H_{t}H_{t}^{\top}=I_{d}$, Lévy’s characterization implies that $\overline{B}^{\,\mathrm{ref}}$ is again a Brownian motion. After $\tau_{c}$, both copies evolve synchronously and therefore stick together.
Lemma 10 (Radial dynamics under reflection coupling).
Fix $s_{0}\in\mathcal{S}_{\mathrm{adm}}$, and let $Y,\bar{Y}$ be coupled as in Definition 5 on $t_{1},t_{2}]\subseteq[0,t_{s}]$. Then there exists a one-dimensional Brownian motion $W$ such that, for $t<\tau_{c}$,
$\mathop{}\!\mathrm{d}r_{t}=\Big\langle\frac{D_{t}}{\|D_{t}\|},\,\widehat{b}_{t}(Y_{t})-\widehat{b}_{t}(\bar{Y}_{t})\Big\rangle\,\mathop{}\!\mathrm{d}t+2g(T-t)\,\mathop{}\!\mathrm{d}W_{t}.$ (31)
Moreover,
$\Big\langle\frac{D_{t}}{\|D_{t}\|},\,\widehat{b}_{t}(Y_{t})-\widehat{b}_{t}(\bar{Y}_{t})\Big\rangle\leq-\kappa_{\widehat{b}_{t}}(r_{t})\,r_{t},$ (32)
and hence
$\mathop{}\!\mathrm{d}r_{t}\leq-\underline{\kappa}_{s_{0}}(r_{t})\,r_{t}\,\mathop{}\!\mathrm{d}t+2g(T-t)\,\mathop{}\!\mathrm{d}W_{t},\qquad t<\tau_{c}.$ (33)
Proof.
For $t<\tau_{c}$, one has $n_{t}=D_{t}/r_{t}$ and
$I_{d}-H_{t}=2n_{t}n_{t}^{\top}.$
Hence
$\mathop{}\!\mathrm{d}D_{t}=\bigl(\widehat{b}_{t}(Y_{t})-\widehat{b}_{t}(\bar{Y}_{t})\bigr)\,\mathop{}\!\mathrm{d}t+2g(T-t)\,n_{t}n_{t}^{\top}\,\mathop{}\!\mathrm{d}\overline{B}_{t}=\bigl(\widehat{b}_{t}(Y_{t})-\widehat{b}_{t}(\bar{Y}_{t})\bigr)\,\mathop{}\!\mathrm{d}t+2g(T-t)\,n_{t}\,\mathop{}\!\mathrm{d}W_{t},$
where
$W_{t}-W_{t_{1}}:=\int_{t_{1}}^{t}\langle n_{u},\mathop{}\!\mathrm{d}\overline{B}_{u}\rangle.$
Since $n_{t}$ is predictable and $\|n_{t}\|=1$ on $\{t<\tau_{c}\}$, $W$ is a one-dimensional Brownian motion.
To justify the radial equation, localize away from the origin. For $n\geq 1$, let
$\tau_{n}:=\inf\{t\geq t_{1}:\ r_{t}\leq n^{-1}\}.$
On $t_{1},\tau_{n}]$, the map $x\mapsto\|x\|$ is $C^{2}$ in a neighborhood of $D_{t}$, so Itô’s formula gives
$\mathop{}\!\mathrm{d}r_{t}=\left\langle\frac{D_{t}}{\|D_{t}\|},\,\widehat{b}_{t}(Y_{t})-\widehat{b}_{t}(\bar{Y}_{t})\right\rangle\mathop{}\!\mathrm{d}t+2g(T-t)\,\mathop{}\!\mathrm{d}W_{t}+\frac{1}{2}\operatorname{Tr}\!\bigl(4g(T-t)^{2}n_{t}n_{t}^{\top}\nabla^{2}\|D_{t}\|\bigr)\,\mathop{}\!\mathrm{d}t.$
Now
$\nabla^{2}\|x\|=\frac{1}{\|x\|}\Big(I_{d}-\frac{xx^{\top}}{\|x\|^{2}}\Big),\qquad x\neq 0,$
and $n_{t}=D_{t}/\|D_{t}\|$, so
$\operatorname{Tr}\!\bigl(n_{t}n_{t}^{\top}\nabla^{2}\|D_{t}\|\bigr)=0.$
Thus the second-order term vanishes identically. Letting $n\to\infty$ proves (31) for $t<\tau_{c}$.
The bound (32) is immediate from Definition 2, because $r_{t}=\|D_{t}\|$. Combining it with Lemma 9 gives (33). ∎
C.3 Construction and generator bound for the switch metric
Definition 6 (Radial-cost Wasserstein metric).
Let $\varphi:[0,\infty)\to[0,\infty)$ be increasing, concave, and satisfy $\varphi(0)=0$. For $\mu,\nu\in\mathcal{P}_{1}(\mathbb{R}^{d})$, define
$W_{\varphi}(\mu,\nu):=\inf_{\pi\in\Pi(\mu,\nu)}\int_{\mathbb{R}^{d}\times\mathbb{R}^{d}}\varphi(\|x-y\|)\,\pi(\mathop{}\!\mathrm{d}x,\mathop{}\!\mathrm{d}y).$
For such $\varphi$, the concavity and monotonicity of $\varphi$ imply subadditivity, and the usual gluing argument shows that $W_{\varphi}$ is a transport metric on $\mathcal{P}_{1}(\mathbb{R}^{d})$.
The remaining task is to choose a concave cost whose generator absorbs the two-zone radial inequality from Lemma 9. On the core $0,R_{\mathrm{sw}}(s_{0})]$, the drift may still expand, so one needs negative curvature of $\varphi$ to be paid for by the diffusion term. Beyond $R_{\mathrm{sw}}(s_{0})$, the drift is already uniformly contractive, so further concavity would only waste this reserve. This leads to a Gaussian core and an affine tail.
For the fixed switch $s_{0}$, define
$\lambda(s_{0}):=\frac{\overline{\mathfrak{b}}(s_{0})^{+}+m_{\mathrm{sw}}(s_{0})}{4\,\underline{g}(s_{0})^{2}},\qquad\mathfrak{a}(s_{0}):=e^{-\lambda(s_{0})R_{\mathrm{sw}}(s_{0})^{2}},\qquad c(s_{0}):=m_{\mathrm{sw}}(s_{0})\,\mathfrak{a}(s_{0}).$ (34)
Attach to $s_{0}$ the cost
$\varphi_{s_{0}}(r):=\begin{cases}\displaystyle\int_{0}^{r}e^{-\lambda(s_{0})u^{2}}\,\mathop{}\!\mathrm{d}u,&0\leq r\leq R_{\mathrm{sw}}(s_{0}),\\[8.00003pt] \displaystyle\varphi_{s_{0}}(R_{\mathrm{sw}}(s_{0}))+\mathfrak{a}(s_{0})\bigl(r-R_{\mathrm{sw}}(s_{0})\bigr),&r\geq R_{\mathrm{sw}}(s_{0}).\end{cases}$ (35)
Lemma 11 (The switch gauge and its generator inequality).
Fix $s_{0}\in\mathcal{S}_{\mathrm{adm}}$ such that $\underline{g}(s_{0})>0$, $\overline{\mathfrak{b}}(s_{0})<\infty$, and $\overline{G}(s_{0})<\infty$. Then $\varphi_{s_{0}}\in C^{1}([0,\infty))$ is increasing and concave, and for all $r\geq 0$,
$\mathfrak{a}(s_{0})\,r\leq\varphi_{s_{0}}(r)\leq r.$ (36)
Moreover,
$2\underline{g}(s_{0})^{2}\varphi_{s_{0}}^{\prime\prime}(r)-\underline{\kappa}_{s_{0}}(r)\,r\,\varphi_{s_{0}}^{\prime}(r)\leq-c(s_{0})\,\varphi_{s_{0}}(r)\qquad\text{for a.e. }r>0.$ (37)
Consequently, $W_{\varphi_{s_{0}}}$ is finite on $\mathcal{P}_{1}(\mathbb{R}^{d})$ and satisfies the triangle inequality.
Proof.
For brevity write
$\lambda:=\lambda(s_{0}),\quad R:=R_{\mathrm{sw}}(s_{0}),\quad\underline{g}:=\underline{g}(s_{0}),\quad\overline{\mathfrak{b}}:=\overline{\mathfrak{b}}(s_{0}),\quad m:=m_{\mathrm{sw}}(s_{0}),$
$\mathfrak{a}:=\mathfrak{a}(s_{0}),\quad c:=c(s_{0}),\quad\underline{\kappa}:=\underline{\kappa}_{s_{0}},\quad\varphi:=\varphi_{s_{0}}.$
From the definition,
$\varphi^{\prime}(r)=\begin{cases}e^{-\lambda r^{2}},&0<r<R,\\ e^{-\lambda R^{2}},&r>R,\end{cases}\qquad\varphi^{\prime\prime}(r)=\begin{cases}-2\lambda r\,e^{-\lambda r^{2}},&0<r<R,\\ 0,&r>R,\end{cases}$
for a.e. $r>0$. Hence $\varphi$ is increasing and concave, and $\varphi^{\prime}$ is continuous at $R$, so $\varphi\in C^{1}([0,\infty))$.
The upper bound $\varphi(r)\leq r$ follows from $0\leq\varphi^{\prime}\leq 1$. For the lower bound, if $0\leq r\leq R$, then
$\varphi(r)=\int_{0}^{r}e^{-\lambda u^{2}}\,\mathop{}\!\mathrm{d}u\geq re^{-\lambda R^{2}}=\mathfrak{a}r.$
If $r\geq R$, then
$\varphi(r)=\varphi(R)+\mathfrak{a}(r-R)\geq\mathfrak{a}R+\mathfrak{a}(r-R)=\mathfrak{a}r.$
This proves (36).
If $0<r\leq R$, then $\underline{\kappa}(r)=-\overline{\mathfrak{b}}$, and therefore
$2\underline{g}^{2}\varphi^{\prime\prime}(r)-\underline{\kappa}(r)r\varphi^{\prime}(r)=\bigl(\overline{\mathfrak{b}}-4\underline{g}^{2}\lambda\bigr)re^{-\lambda r^{2}}.$
By the definition of $\lambda$,
$\overline{\mathfrak{b}}-4\underline{g}^{2}\lambda=\overline{\mathfrak{b}}-\overline{\mathfrak{b}}^{+}-m\leq-m.$
Hence
$2\underline{g}^{2}\varphi^{\prime\prime}(r)-\underline{\kappa}(r)r\varphi^{\prime}(r)\leq-mre^{-\lambda r^{2}}.$
Since $r\leq R$ implies $e^{-\lambda r^{2}}\geq e^{-\lambda R^{2}}=\mathfrak{a}$, and $\varphi(r)\leq r$,
$c\,\varphi(r)=m\mathfrak{a}\,\varphi(r)\leq me^{-\lambda r^{2}}r.$
Thus the generator inequality holds on $(0,R]$.
If $r>R$, then $\underline{\kappa}(r)=m$, $\varphi^{\prime\prime}(r)=0$, and $\varphi^{\prime}(r)=\mathfrak{a}$, so
$2\underline{g}^{2}\varphi^{\prime\prime}(r)-\underline{\kappa}(r)r\varphi^{\prime}(r)=-m\mathfrak{a}\,r=-cr\leq-c\,\varphi(r),$
again because $\varphi(r)\leq r$. This proves (37). ∎
C.4 Early semigroup contraction
Remark 6 (The extra theorem-level input used in the early phase).
Beyond Assumptions 12, the next proposition uses one additional theorem-level input: the learned reverse SDE must be strongly well posed on $0,t_{s}]$ and must admit coalescing reflection couplings on each subinterval for every initial coupling. This is not automatic from the standing assumptions. Later, Proposition 7 and Corollary 4 give convenient sufficient conditions, including the standard global/local Lipschitz regimes used in the literature.
Lemma 12 (Signed-load Grönwall for synchronized differences).
Let $D:[u,v]\to\mathbb{R}^{d}$ be absolutely continuous. Assume that $\beta\in L^{1}(u,v)$, $F\in L^{1}(u,v;\mathbb{R}^{d})$, and for a.e. $t\in[u,v]$,
$\dot{D}_{t}=A_{t}+F_{t},\qquad\langle D_{t},A_{t}\rangle\leq\beta(t)\|D_{t}\|^{2}.$
Then for every $t\in[u,v]$,
$\|D_{t}\|\leq e^{\int_{u}^{t}\beta(r)\,\mathop{}\!\mathrm{d}r}\|D_{u}\|+\int_{u}^{t}e^{\int_{r}^{t}\beta(\tau)\,\mathop{}\!\mathrm{d}\tau}\|F_{r}\|\,\mathop{}\!\mathrm{d}r.$
Proof.
For $\varepsilon>0$, define
$q_{\varepsilon}(t):=\sqrt{\|D_{t}\|^{2}+\varepsilon}.$
Since $D$ is absolutely continuous, so is $q_{\varepsilon}$, and for a.e. $t$,
$q_{\varepsilon}^{\prime}(t)=\frac{\langle D_{t},A_{t}+F_{t}\rangle}{q_{\varepsilon}(t)}\leq\beta(t)\frac{\|D_{t}\|^{2}}{q_{\varepsilon}(t)}+\|F_{t}\|.$
Now
$\beta(t)\frac{\|D_{t}\|^{2}}{q_{\varepsilon}(t)}=\beta(t)q_{\varepsilon}(t)-\beta(t)\frac{\varepsilon}{q_{\varepsilon}(t)}\leq\beta(t)q_{\varepsilon}(t)+|\beta(t)|\sqrt{\varepsilon},$
because $q_{\varepsilon}(t)\geq\sqrt{\varepsilon}$. Hence
$q_{\varepsilon}^{\prime}(t)\leq\beta(t)q_{\varepsilon}(t)+\|F_{t}\|+|\beta(t)|\sqrt{\varepsilon}.$
Gronwall’s inequality yields
$q_{\varepsilon}(t)\leq e^{\int_{u}^{t}\beta(\tau)\,\mathop{}\!\mathrm{d}\tau}\,q_{\varepsilon}(u)+\int_{u}^{t}e^{\int_{r}^{t}\beta(\tau)\,\mathop{}\!\mathrm{d}\tau}\|F_{r}\|\,\mathop{}\!\mathrm{d}r+\sqrt{\varepsilon}\int_{u}^{t}e^{\int_{r}^{t}\beta(\tau)\,\mathop{}\!\mathrm{d}\tau}|\beta(r)|\,\mathop{}\!\mathrm{d}r.$
Letting $\varepsilon\downarrow 0$ proves the claim. ∎
Lemma 13.
Fix $s_{0}\in\mathcal{S}_{\mathrm{adm}}$ such that $\underline{g}(s_{0})>0$, $\overline{\mathfrak{b}}(s_{0})<\infty$, and $\overline{G}(s_{0})<\infty$. Assume, in addition, that the learned reverse SDE is strongly well posed on $0,t_{s}]$ and admits coalescing reflection couplings on every subinterval $u,v]\subseteq[0,t_{s}]$ for every initial coupling $\pi\in\Pi(\mu,\nu)$ with $\mu,\nu\in\mathcal{P}_{1}(\mathbb{R}^{d})$. Then
$W_{\varphi_{s_{0}}}\!\bigl(\mu P_{u,v},\nu P_{u,v}\bigr)\leq e^{-c(s_{0})(v-u)}\,W_{\varphi_{s_{0}}}(\mu,\nu),\qquad 0\leq u\leq v\leq t_{s},$ (38)
for every $\mu,\nu\in\mathcal{P}_{1}(\mathbb{R}^{d})$.
Proof.
Fix $\mu,\nu\in\mathcal{P}_{1}(\mathbb{R}^{d})$ and $0\leq u\leq v\leq t_{s}$. Choose any coupling $\pi\in\Pi(\mu,\nu)$, and let $(Y_{t},\bar{Y}_{t})$ be a coalescing reflection coupling of the learned reverse SDE on $[u,v]$ in the sense of Definition 5, with initial law $\pi$. Write
$r_{t}:=\|Y_{t}-\bar{Y}_{t}\|,\qquad\varphi:=\varphi_{s_{0}},\qquad c:=c(s_{0}),\qquad\underline{g}:=\underline{g}(s_{0}).$
By Lemma 10,
$\mathop{}\!\mathrm{d}r_{t}\leq-\underline{\kappa}_{s_{0}}(r_{t})\,r_{t}\,\mathop{}\!\mathrm{d}t+2g(T-t)\,\mathop{}\!\mathrm{d}W_{t}\qquad\text{for }t<\tau_{c}.$
Apply Meyer–Itô’s formula to the stopped process $t\mapsto\varphi(r_{t\wedge\tau_{c}})$. This is legitimate because $\varphi\in C^{1}$ and is piecewise $C^{2}$, and the only potential local-time term at the gluing point $R_{\mathrm{sw}}(s_{0})$ vanishes because $\varphi^{\prime}$ is continuous there. Using $\varphi^{\prime\prime}\leq 0$ and (33), we obtain
$\mathop{}\!\mathrm{d}\varphi(r_{t\wedge\tau_{c}})\leq\Bigl(2g(T-t)^{2}\varphi^{\prime\prime}(r_{t})-\underline{\kappa}_{s_{0}}(r_{t})\,r_{t}\,\varphi^{\prime}(r_{t})\Bigr)\mathbf{1}_{\{t<\tau_{c}\}}\mathop{}\!\mathrm{d}t+2g(T-t)\varphi^{\prime}(r_{t})\mathbf{1}_{\{t<\tau_{c}\}}\mathop{}\!\mathrm{d}W_{t}.$
Since $g(T-t)\geq\underline{g}$ on $[u,v]$ and $\varphi^{\prime\prime}\leq 0$,
$2g(T-t)^{2}\varphi^{\prime\prime}(r_{t})\leq 2\underline{g}^{2}\varphi^{\prime\prime}(r_{t}).$
Hence Lemma 11 gives
$\mathop{}\!\mathrm{d}\varphi(r_{t\wedge\tau_{c}})\leq-c\,\varphi(r_{t\wedge\tau_{c}})\,\mathop{}\!\mathrm{d}t+2g(T-t)\varphi^{\prime}(r_{t})\mathbf{1}_{\{t<\tau_{c}\}}\mathop{}\!\mathrm{d}W_{t}.$
Because $0\leq\varphi^{\prime}\leq 1$ and $\int_{u}^{v}g(T-t)^{2}\,\mathop{}\!\mathrm{d}t<\infty$, the stochastic integral is square integrable and therefore a true martingale. Multiplying by $e^{ct}$, integrating from $u$ to $v$, and taking expectations yields
$\mathop{\mathbb{E}}\nolimits[\varphi(r_{v})]\leq e^{-c(v-u)}\,\mathop{\mathbb{E}}\nolimits[\varphi(r_{u})].$
Since the law of $(Y_{v},\bar{Y}_{v})$ is a coupling of $\mu P_{u,v}$ and $\nu P_{u,v}$,
$W_{\varphi}(\mu P_{u,v},\nu P_{u,v})\leq\mathop{\mathbb{E}}\nolimits[\varphi(r_{v})]\leq e^{-c(v-u)}\,\mathop{\mathbb{E}}\nolimits[\varphi(r_{u})].$
Taking the infimum over $\pi\in\Pi(\mu,\nu)$ proves (38). ∎
Once semigroup contraction is available, every pre-switch error source is handled in the same way: insert the error at the time it enters the chain and propagate it to $t_{s}$ with the damping factor $e^{-c(s_{0})(t_{s}-\cdot)}$. The next proposition separates the three places where error enters: initialization, discretization, and score forcing.
Remark 7 (Why the coupling changes).
For the learned reverse SDE, synchronous coupling only sees the near-field endpoint $\mathfrak{b}(s)$, because formally
$\frac{1}{2}\frac{\mathop{}\!\mathrm{d}}{\mathop{}\!\mathrm{d}t}\|Y_{t}-\bar{Y}_{t}\|^{2}\leq\mathfrak{b}(s)\|Y_{t}-\bar{Y}_{t}\|^{2}.$
Reflection coupling instead acts on the radial process and therefore accesses the full profile $r\mapsto\kappa_{\widehat{b}_{t}}(r)$, including its contractive tail. This is the only reason the proof switches couplings.
Lemma 14.
Fix $s_{0}\in\mathcal{S}_{\mathrm{adm}}$ with $t_{K}=t_{s}:=T-s_{0}$, and assume $\widehat{p}_{T}$ belongs to $\mathcal{P}_{1}(\mathbb{R}^{d})$, Assumption 2, the hypotheses of Lemma 13, and that the theorem-level ideal reverse SDE (12) admits a global strong solution
$(\widetilde{X}_{t})_{0\leq t\leq T}$
with
$\operatorname{Law}(\widetilde{X}_{t})=p_{T-t},\qquad 0\leq t\leq T.$
Then
$\Delta_{\varphi}^{\mathrm{sw}}(s_{0}):=W_{\varphi_{s_{0}}}\!\bigl(\operatorname{Law}(Z_{t_{s}}),p_{s_{0}}\bigr)$
satisfies
$\Delta_{\varphi}^{\mathrm{sw}}(s_{0})\leq\;e^{-c(s_{0})t_{s}}W_{\varphi_{s_{0}}}(\widehat{p}_{T},p_{T})$ $+\sum_{k=0}^{K-1}e^{-c(s_{0})(t_{s}-t_{k+1})}\Bigl(\mathop{\mathbb{E}}\nolimits[\xi_{k}(Z_{t_{k}})^{2}]\Bigr)^{1/2}$ (39)
$+\int_{s_{0}}^{T}e^{-c(s_{0})(s-s_{0})}\,g^{2}(s)\,\varepsilon(s)\,\mathop{}\!\mathrm{d}s.$
In particular, the right-hand side is exactly $\overline{\Delta}_{\varphi}^{\mathrm{sw}}(s_{0})$ from (27).
Proof.
Let $\varphi:=\varphi_{s_{0}}$ and $c:=c(s_{0})$. Write
$\mu_{k}:=\operatorname{Law}(Z_{t_{k}}),\qquad\rho_{k}:=\mu_{k}P_{t_{k},t_{s}},\qquad k=0,\dots,K.$
Then
$\rho_{0}=\widehat{p}_{T}P_{0,t_{s}},\qquad\rho_{K}=\operatorname{Law}(Z_{t_{s}}).$
Initialization mismatch.
By Lemma 13,
$W_{\varphi}\!\bigl(\widehat{p}_{T}P_{0,t_{s}},\,p_{T}P_{0,t_{s}}\bigr)\leq e^{-ct_{s}}W_{\varphi}(\widehat{p}_{T},p_{T}).$
Early discretization.
By the triangle inequality for $W_{\varphi}$,
$W_{\varphi}(\rho_{K},\rho_{0})\leq\sum_{k=0}^{K-1}W_{\varphi}(\rho_{k+1},\rho_{k}).$
Since
$\rho_{k+1}=(\mu_{k}\Psi_{k})P_{t_{k+1},t_{s}},\qquad\rho_{k}=(\mu_{k}P_{k})P_{t_{k+1},t_{s}},$
Lemma 13 gives
$W_{\varphi}(\rho_{k+1},\rho_{k})\leq e^{-c(t_{s}-t_{k+1})}\,W_{\varphi}(\mu_{k}\Psi_{k},\mu_{k}P_{k}).$
Now $\varphi(r)\leq r$, so $W_{\varphi}\leq W_{1}\leq W_{2}$. Apply Zhang [2013, Theorem 1.1] on the parameter space $(\mathbb{R}^{d},\mathcal{B}(\mathbb{R}^{d}),\mu_{k})$ with
$\mu_{x}:=\Psi_{k}(x,\cdot),\qquad\nu_{x}:=P_{k}(x,\cdot),\qquad c(y,z):=\|y-z\|^{2}.$
This yields a Borel measurable family
$x\longmapsto\pi_{k}^{x}\in\Pi(\Psi_{k}(x,\cdot),P_{k}(x,\cdot))$
of optimal couplings for the quadratic cost. Integrating $\pi_{k}^{x}$ against $\mu_{k}(\mathop{}\!\mathrm{d}x)$ gives a coupling of $\mu_{k}\Psi_{k}$ and $\mu_{k}P_{k}$, and therefore
$W_{2}^{2}(\mu_{k}\Psi_{k},\mu_{k}P_{k})\leq\int_{\mathbb{R}^{d}}W_{2}^{2}\!\bigl(\Psi_{k}(x,\cdot),P_{k}(x,\cdot)\bigr)\,\mu_{k}(\mathop{}\!\mathrm{d}x)\leq\mathop{\mathbb{E}}\nolimits[\xi_{k}(Z_{t_{k}})^{2}].$
Hence
$W_{\varphi}(\mu_{k}\Psi_{k},\mu_{k}P_{k})\leq W_{2}(\mu_{k}\Psi_{k},\mu_{k}P_{k})\leq\Bigl(\mathop{\mathbb{E}}\nolimits[\xi_{k}(Z_{t_{k}})^{2}]\Bigr)^{1/2}.$
Summing over $k$ gives
$W_{\varphi}\!\bigl(\operatorname{Law}(Z_{t_{s}}),\widehat{p}_{T}P_{0,t_{s}}\bigr)\leq\sum_{k=0}^{K-1}e^{-c(t_{s}-t_{k+1})}\Bigl(\mathop{\mathbb{E}}\nolimits[\xi_{k}(Z_{t_{k}})^{2}]\Bigr)^{1/2}.$ (40)
Pre-switch score forcing.
This step uses two different couplings. To propagate a law discrepancy from time $u_{j+1}$ to $t_{s}$, we use the learned-semigroup contraction from Lemma 13, whose proof is based on reflection coupling. But to compare the ideal and learned reverse dynamics on the short interval $u_{j},u_{j+1}]$, we use a synchronous comparison built from the global ideal reverse path. No separate restart assumption for the ideal equation is used: we simply restrict the theorem-level solution $(\widetilde{X}_{t})_{0\leq t\leq T}$ of (12) to the interval $u_{j},u_{j+1}]$, and solve the learned equation on the same filtered space and with the same Brownian motion $\overline{B}$, started from the random state $\widetilde{X}_{u_{j}}$.
Let $0=u_{0}<\cdots<u_{m}=t_{s}$ be a partition, and define
$\eta_{j}:=p_{T-u_{j}}P_{u_{j},t_{s}},\qquad j=0,\dots,m.$
Then $\eta_{0}=p_{T}P_{0,t_{s}}$ and $\eta_{m}=p_{s_{0}}$, so
$W_{\varphi}(p_{T}P_{0,t_{s}},p_{s_{0}})\leq\sum_{j=0}^{m-1}W_{\varphi}(\eta_{j},\eta_{j+1}).$
Fix $j$. Let $Y^{(j)}$ be the strong solution of the learned reverse SDE on $u_{j},u_{j+1}]$, driven by the same Brownian motion $\overline{B}$ as $\widetilde{X}$, and initialized by
$Y^{(j)}_{u_{j}}=\widetilde{X}_{u_{j}}.$
Since $\operatorname{Law}(\widetilde{X}_{u_{j}})=p_{T-u_{j}}$, the Markov property of the learned reverse semigroup gives
$\operatorname{Law}\!\bigl(Y^{(j)}_{u_{j+1}}\mid\mathcal{F}_{u_{j}}\bigr)=P_{u_{j},u_{j+1}}(\widetilde{X}_{u_{j}},\cdot),$
and therefore
$\operatorname{Law}\!\bigl(Y^{(j)}_{u_{j+1}}\bigr)=p_{T-u_{j}}P_{u_{j},u_{j+1}}.$
On the other hand, because $\widetilde{X}$ is the global ideal reverse trajectory,
$\operatorname{Law}(\widetilde{X}_{u_{j+1}})=p_{T-u_{j+1}}.$
Therefore
$\eta_{j}=\operatorname{Law}\!\bigl(Y^{(j)}_{u_{j+1}}\bigr)P_{u_{j+1},t_{s}},\qquad\eta_{j+1}=\operatorname{Law}(\widetilde{X}_{u_{j+1}})P_{u_{j+1},t_{s}}.$
Applying Lemma 13 from $u_{j+1}$ to $t_{s}$ gives
$W_{\varphi}(\eta_{j},\eta_{j+1})\leq e^{-c(t_{s}-u_{j+1})}\,W_{\varphi}\!\bigl(\operatorname{Law}(Y^{(j)}_{u_{j+1}}),\operatorname{Law}(\widetilde{X}_{u_{j+1}})\bigr).$
Set $D_{t}:=Y^{(j)}_{t}-\widetilde{X}_{t}$. Because the Brownian motions are identical, the noise cancels and
$\dot{D}_{t}=\widehat{b}_{t}(Y^{(j)}_{t})-\widehat{b}_{t}(\widetilde{X}_{t})+\delta_{t}(\widetilde{X}_{t})\qquad\text{for a.e. }t\in[u_{j},u_{j+1}].$
By Lemma 1(iii),
$\langle D_{t},\widehat{b}_{t}(Y^{(j)}_{t})-\widehat{b}_{t}(\widetilde{X}_{t})\rangle\leq\mathfrak{b}(T-t)\|D_{t}\|^{2}\leq\overline{\mathfrak{b}}(s_{0})\|D_{t}\|^{2}$
for a.e. $t\in[u_{j},u_{j+1}]$. Lemma 12, with
$A_{t}:=\widehat{b}_{t}(Y^{(j)}_{t})-\widehat{b}_{t}(\widetilde{X}_{t}),\qquad F_{t}:=\delta_{t}(\widetilde{X}_{t}),\qquad\beta(t):=\overline{\mathfrak{b}}(s_{0}),$
and $D_{u_{j}}=0$, therefore yields
$\|D_{u_{j+1}}\|\leq\int_{u_{j}}^{u_{j+1}}e^{\overline{\mathfrak{b}}(s_{0})(u_{j+1}-r)}\|\delta_{r}(\widetilde{X}_{r})\|\,\mathop{}\!\mathrm{d}r.$
Since $W_{\varphi}\leq W_{1}$, the coupling $(Y^{(j)}_{u_{j+1}},\widetilde{X}_{u_{j+1}})$ gives
$W_{\varphi}\!\bigl(\operatorname{Law}(Y^{(j)}_{u_{j+1}}),\operatorname{Law}(\widetilde{X}_{u_{j+1}})\bigr)\leq\mathop{\mathbb{E}}\nolimits\|D_{u_{j+1}}\|\leq\int_{u_{j}}^{u_{j+1}}e^{\overline{\mathfrak{b}}(s_{0})(u_{j+1}-r)}\mathop{\mathbb{E}}\nolimits\|\delta_{r}(\widetilde{X}_{r})\|\,\mathop{}\!\mathrm{d}r.$
Therefore
$W_{\varphi}(p_{T}P_{0,t_{s}},p_{s_{0}})\leq\sum_{j=0}^{m-1}\int_{u_{j}}^{u_{j+1}}e^{-c(t_{s}-u_{j+1})}e^{\overline{\mathfrak{b}}(s_{0})(u_{j+1}-r)}\mathop{\mathbb{E}}\nolimits\|\delta_{r}(\widetilde{X}_{r})\|\,\mathop{}\!\mathrm{d}r.$
Now $\widetilde{X}_{r}\sim p_{T-r}$ for each $r\in[0,t_{s}]$ by the theorem-level hypothesis, and
$\mathop{\mathbb{E}}\nolimits\|\delta_{r}(\widetilde{X}_{r})\|=g^{2}(T-r)\,\mathop{\mathbb{E}}\nolimits_{X\sim p_{T-r}}\|e_{T-r}(X)\|\leq g^{2}(T-r)\varepsilon(T-r)$
by Cauchy–Schwarz and Assumption 2. Hence the integrand is dominated by
$e^{(\overline{\mathfrak{b}}(s_{0})+c)t_{s}}\,g^{2}(T-r)\varepsilon(T-r),$
which is integrable on $0,t_{s}]$. Letting the mesh of the partition tend to zero and using dominated convergence gives
$W_{\varphi}(p_{T}P_{0,t_{s}},p_{s_{0}})\leq\int_{0}^{t_{s}}e^{-c(t_{s}-r)}\,\mathop{\mathbb{E}}\nolimits\|\delta_{r}(\widetilde{X}_{r})\|\,\mathop{}\!\mathrm{d}r=\int_{s_{0}}^{T}e^{-c(s-s_{0})}\,g^{2}(s)\,\mathop{\mathbb{E}}\nolimits_{X\sim p_{s}}\|e_{s}(X)\|\,\mathop{}\!\mathrm{d}s.$
Applying Cauchy–Schwarz once more yields
$W_{\varphi}(p_{T}P_{0,t_{s}},p_{s_{0}})\leq\int_{s_{0}}^{T}e^{-c(s-s_{0})}\,g^{2}(s)\,\varepsilon(s)\,\mathop{}\!\mathrm{d}s.$ (41)
Combine the three pieces.
By the triangle inequality,
$W_{\varphi}(\operatorname{Law}(Z_{t_{s}}),p_{s_{0}})\leq W_{\varphi}(\operatorname{Law}(Z_{t_{s}}),\widehat{p}_{T}P_{0,t_{s}})+W_{\varphi}(\widehat{p}_{T}P_{0,t_{s}},p_{T}P_{0,t_{s}})+W_{\varphi}(p_{T}P_{0,t_{s}},p_{s_{0}}).$
Substituting (40), the initialization estimate, and (41) proves (39). ∎
Appendix D Switch-time conversion back to $W_{2}$
This section performs the one-time return to Euclidean transport used by our one-switch family. The proof is deliberately specialized to the present switch metric. Because $\varphi_{s_{0}}$ has an affine tail, one can decompose $D^{2}$ into a small-distance piece controlled by $\varphi_{s_{0}}(D)$ and a large-distance piece paid for by a $p$-moment. This produces the exponent $\theta_{p}$ without any additional interpolation machinery and explains why the later sharpness statement is tied to affine-tail costs.
Lemma 15.
Fix an admissible switch $s_{0}\in\mathcal{S}_{\mathrm{adm}}$ and $p>2$. Assume that the switch-time budget $\overline{\mathcal{M}}_{p}^{\mathrm{sw}}(s_{0})$ from (25) is finite. Then
$W_{2}\!\bigl(\operatorname{Law}(Z_{t_{s}}),p_{s_{0}}\bigr)\leq C_{p}^{\mathrm{sw}}(s_{0})\,\bigl(\Delta_{\varphi}^{\mathrm{sw}}(s_{0})\bigr)^{\theta_{p}},$ (42)
where $C_{p}^{\mathrm{sw}}(s_{0})$ is the switch factor from Section 4.
Proof.
The argument is a one-threshold decomposition of $D^{2}$. Small separations are controlled by the linear lower bound $\varphi(r)\geq\mathfrak{a}r$, while large separations are paid for by the $p$-moment budget. Let $\varphi:=\varphi_{s_{0}}$, $\mathfrak{a}:=\mathfrak{a}(s_{0})$, and
$\mu:=\operatorname{Law}(Z_{t_{s}}),\qquad\nu:=p_{s_{0}}.$
By Lemma 11, $\varphi(r)\geq\mathfrak{a}r$ for all $r\geq 0$.
Fix any coupling $(U,V)$ of $(\mu,\nu)$, and set $D:=\|U-V\|$. For every $\rho>0$,
$D^{2}=D^{2}\mathbf{1}_{\{D\leq\rho\}}+D^{2}\mathbf{1}_{\{D>\rho\}}\leq\rho D+\rho^{-(p-2)}D^{p}\leq\frac{\rho}{\mathfrak{a}}\,\varphi(D)+\rho^{-(p-2)}D^{p}.$
Taking expectations and using
$\mathop{\mathbb{E}}\nolimits[D^{p}]\leq 2^{p-1}\bigl(\mathop{\mathbb{E}}\nolimits\|U\|^{p}+\mathop{\mathbb{E}}\nolimits\|V\|^{p}\bigr)\leq 2^{p-1}\,\overline{\mathcal{M}}_{p}^{\mathrm{sw}}(s_{0})$
gives
$\mathop{\mathbb{E}}\nolimits[D^{2}]\leq\frac{\rho}{\mathfrak{a}}\,\mathop{\mathbb{E}}\nolimits[\varphi(D)]+\frac{2^{p-1}\overline{\mathcal{M}}_{p}^{\mathrm{sw}}(s_{0})}{\rho^{p-2}}.$
Taking the infimum over couplings yields
$W_{2}^{2}(\mu,\nu)\leq\frac{\rho}{\mathfrak{a}}\,W_{\varphi}(\mu,\nu)+\frac{2^{p-1}\overline{\mathcal{M}}_{p}^{\mathrm{sw}}(s_{0})}{\rho^{p-2}}.$
Optimizing in $\rho$ gives
$W_{2}(\mu,\nu)\leq C_{\mathrm{tr}}(p)\,\mathfrak{a}(s_{0})^{-\theta_{p}}\,\bigl(\overline{\mathcal{M}}_{p}^{\mathrm{sw}}(s_{0})\bigr)^{\frac{1}{2(p-1)}}\,\bigl(W_{\varphi}(\mu,\nu)\bigr)^{\theta_{p}},$
which is exactly (42). ∎
D.1 An explicit admissible moment criterion
Proposition 2.
Fix $p>2$, an admissible switch $s_{0}\in\mathcal{S}_{\mathrm{adm}}$, and let $t_{s}:=T-s_{0}$. Assume the grid contains $t_{s}$, say $t_{K}=t_{s}$, and suppose
$\int_{\mathbb{R}^{d}}\|y\|^{p}\,\Psi_{k}(x,\mathop{}\!\mathrm{d}y)\leq(1+A_{k})\|x\|^{p}+B_{k},\qquad x\in\mathbb{R}^{d},\quad k=0,\dots,K-1,$
for constants $A_{k},B_{k}\geq 0$. Then
$\mathop{\mathbb{E}}\nolimits\|Z_{t_{s}}\|^{p}\leq M_{p,\mathrm{num}}^{\mathrm{sw}}(s_{0}):=\Bigg(\prod_{j=0}^{K-1}(1+A_{j})\Bigg)\mathop{\mathbb{E}}\nolimits_{Z_{0}\sim\widehat{p}_{T}}\|Z_{0}\|^{p}+\sum_{i=0}^{K-1}B_{i}\prod_{j=i+1}^{K-1}(1+A_{j}).$
Consequently, one admissible choice in (25) is
$\overline{\mathcal{M}}_{p}^{\mathrm{sw}}(s_{0})=M_{p,\mathrm{num}}^{\mathrm{sw}}(s_{0})+\mathop{\mathbb{E}}\nolimits_{X\sim p_{s_{0}}}\|X\|^{p}.$
Proof.
Set $m_{k}:=\mathop{\mathbb{E}}\nolimits\|Z_{t_{k}}\|^{p}$. Since $\operatorname{Law}(Z_{t_{k+1}})=\operatorname{Law}(Z_{t_{k}})\Psi_{k}$, the one-step bound gives
$m_{k+1}\leq(1+A_{k})m_{k}+B_{k}.$
Iterating this scalar recursion yields the stated bound for $\mathop{\mathbb{E}}\nolimits\|Z_{t_{s}}\|^{p}$. The second claim follows from $\operatorname{Law}(\widetilde{X}_{t_{s}})=p_{s_{0}}$ and Lemma 8. ∎
D.2 Sharpness of the conversion exponent
The next proposition explains why the exponent $\theta_{p}=(p-2)/(2(p-1))$ should be viewed as structural rather than incidental. The same affine-tail softness that makes the early metric usable is exactly what limits the price of converting back to $W_{2}$.
Proposition 3.
Fix $p>2$. Let $\varphi:[0,\infty)\to[0,\infty)$ be concave, nondecreasing, and affine on $R_{\varphi},\infty)$ with slope $a_{\varphi}>0$. Then the exponent
$\theta_{p}:=\frac{p-2}{2(p-1)}$
in the interpolation
$W_{2}(\mu,\nu)\lesssim\bigl(\mathcal{M}_{p}(\mu,\nu)\bigr)^{\frac{1}{2(p-1)}}\bigl(W_{\varphi}(\mu,\nu)\bigr)^{\theta_{p}}$
cannot in general be improved within the class of affine-tail concave costs and $p$-moment switch assumptions. More precisely, there exist pairs $(\mu_{R},\nu_{R})$ with
$\sup_{R\geq 1}\mathcal{M}_{p}(\mu_{R},\nu_{R})<\infty$
such that
$W_{\varphi}(\mu_{R},\nu_{R})\asymp R^{1-p},\qquad W_{2}(\mu_{R},\nu_{R})\asymp R^{1-p/2}\asymp\bigl(W_{\varphi}(\mu_{R},\nu_{R})\bigr)^{\theta_{p}}.$
Proof.
Let $e_{1}=(1,0,\dots,0)$ and define
$\mu_{R}:=(1-R^{-p})\delta_{0}+R^{-p}\delta_{Re_{1}},\qquad\nu_{R}:=\delta_{0}.$
Then $\sup_{R\geq 1}\mathcal{M}_{p}(\mu_{R},\nu_{R})<\infty$. Since $\nu_{R}$ is a Dirac mass, the unique coupling gives
$W_{2}^{2}(\mu_{R},\nu_{R})=R^{-p}\|Re_{1}\|^{2}=R^{2-p},\qquad W_{\varphi}(\mu_{R},\nu_{R})=R^{-p}\varphi(R).$
Because $\varphi$ is affine with positive slope for large $R$, one has $\varphi(R)\asymp R$, so $W_{\varphi}(\mu_{R},\nu_{R})\asymp R^{1-p}$. Hence
$\bigl(W_{\varphi}(\mu_{R},\nu_{R})\bigr)^{\theta_{p}}\asymp R^{(1-p)\frac{p-2}{2(p-1)}}=R^{1-p/2}=W_{2}(\mu_{R},\nu_{R}).$
Appendix E Late-window Euclidean propagation
After the switch, the radial geometry has already done all of its work. What remains is a purely Euclidean short-horizon propagation problem. Accordingly, this section uses only Euclidean $W_{2}$ estimates on $t_{s},T]$ and keeps separate the three quantities that arrive there: the converted switch-time discrepancy, late discretization, and late score forcing.
Proposition 4.
Fix $s_{0}\in[0,T]$ and assume that the grid contains the switch time
$t_{s}:=T-s_{0}=t_{K}$
for some $K\in\{0,\dots,N\}$. Assume Assumptions 1, 2, and 3. Assume moreover that $\Gamma(s_{0})<\infty$, that the learned reverse SDE is strongly well posed on $t_{s},T]$, that the ideal reverse SDE (12) admits a global strong solution
$(\widetilde{X}_{t})_{0\leq t\leq T}$
with
$\operatorname{Law}(\widetilde{X}_{t})=p_{T-t},\qquad 0\leq t\leq T,$
that the one-step defects satisfy
$W_{2}\!\bigl(P_{k}(x,\cdot),\Psi_{k}(x,\cdot)\bigr)\leq\xi_{k}(x),\qquad x\in\mathbb{R}^{d},\quad k=K,\dots,N-1,$
and that the relevant second moments are finite. Then
$W_{2}\!\bigl(\operatorname{Law}(Z_{t_{N}}),p_{0}\bigr)\leq\Delta_{\mathrm{late}}^{\mathrm{disc}}(s_{0})+e^{\Gamma(s_{0})}W_{2}\!\bigl(\operatorname{Law}(Z_{t_{s}}),p_{s_{0}}\bigr)+\Delta_{\mathrm{late}}^{\mathrm{force}}(s_{0}).$ (43)
Proof.
Let
$\mu_{k}:=\operatorname{Law}(Z_{t_{k}}),\qquad\nu_{k}:=\mu_{k}P_{t_{k},T},\qquad k=K,\dots,N.$
Then
$\nu_{N}=\operatorname{Law}(Z_{t_{N}}),\qquad\nu_{K}=\mu_{K}P_{t_{s},T}.$
late discretization.
By the triangle inequality,
$W_{2}(\nu_{N},\nu_{K})\leq\sum_{k=K}^{N-1}W_{2}(\nu_{k+1},\nu_{k}).$
Since
$\nu_{k+1}=(\mu_{k}\Psi_{k})P_{t_{k+1},T},\qquad\nu_{k}=(\mu_{k}P_{k})P_{t_{k+1},T},$
it remains to propagate one-step defects from $t_{k+1}$ to $T$.
Fix $k\in\{K,\dots,N-1\}$, and let $\alpha,\beta\in\mathcal{P}_{2}(\mathbb{R}^{d})$. Choose any coupling $\pi\in\Pi(\alpha,\beta)$, let $(Y_{t_{k+1}},\bar{Y}_{t_{k+1}})\sim\pi$, and let $Y,\bar{Y}$ solve the learned reverse SDE on $t_{k+1},T]$ driven by the same Brownian motion. Set
$D_{t}:=Y_{t}-\bar{Y}_{t}.$
Because the noise is synchronized, $D$ is absolutely continuous and
$\dot{D}_{t}=\widehat{b}_{t}(Y_{t})-\widehat{b}_{t}(\bar{Y}_{t})\qquad\text{for a.e. }t\in[t_{k+1},T].$
By Lemma 1(iii),
$\langle D_{t},\widehat{b}_{t}(Y_{t})-\widehat{b}_{t}(\bar{Y}_{t})\rangle\leq\mathfrak{b}(T-t)\|D_{t}\|^{2}\qquad\text{for a.e. }t\in[t_{k+1},T].$
Applying Lemma 12 with
$A_{t}:=\widehat{b}_{t}(Y_{t})-\widehat{b}_{t}(\bar{Y}_{t}),\qquad F_{t}:=0,\qquad\beta(t):=\mathfrak{b}(T-t),$
yields
$\|D_{T}\|\leq e^{\int_{t_{k+1}}^{T}\mathfrak{b}(T-r)\,\mathop{}\!\mathrm{d}r}\|D_{t_{k+1}}\|=e^{\Gamma(T-t_{k+1})}\|D_{t_{k+1}}\|.$
Squaring and taking expectations gives
$\mathop{\mathbb{E}}\nolimits\|D_{T}\|^{2}\leq e^{2\Gamma(T-t_{k+1})}\mathop{\mathbb{E}}\nolimits\|D_{t_{k+1}}\|^{2}.$
Since $\pi$ was arbitrary, taking the infimum over all $\pi\in\Pi(\alpha,\beta)$ yields
$W_{2}\!\bigl(\alpha P_{t_{k+1},T},\beta P_{t_{k+1},T}\bigr)\leq e^{\Gamma(T-t_{k+1})}W_{2}(\alpha,\beta).$
Applying this with $\alpha=\mu_{k}\Psi_{k}$ and $\beta=\mu_{k}P_{k}$ gives
$W_{2}(\nu_{k+1},\nu_{k})\leq e^{\Gamma(T-t_{k+1})}\,W_{2}(\mu_{k}\Psi_{k},\mu_{k}P_{k}).$
Apply Zhang [2013, Theorem 1.1] on the parameter space $(\mathbb{R}^{d},\mathcal{B}(\mathbb{R}^{d}),\mu_{k})$ with
$\mu_{x}:=\Psi_{k}(x,\cdot),\qquad\nu_{x}:=P_{k}(x,\cdot),\qquad c(y,z):=\|y-z\|^{2}.$
This yields a Borel measurable family
$x\longmapsto\pi_{k}^{x}\in\Pi(\Psi_{k}(x,\cdot),P_{k}(x,\cdot))$
of optimal couplings for the quadratic cost. Integrating $\pi_{k}^{x}$ against $\mu_{k}(\mathop{}\!\mathrm{d}x)$ gives a coupling of $\mu_{k}\Psi_{k}$ and $\mu_{k}P_{k}$, and therefore
$W_{2}^{2}(\mu_{k}\Psi_{k},\mu_{k}P_{k})\leq\int_{\mathbb{R}^{d}}W_{2}^{2}\!\bigl(\Psi_{k}(x,\cdot),P_{k}(x,\cdot)\bigr)\,\mu_{k}(\mathop{}\!\mathrm{d}x)\leq\mathop{\mathbb{E}}\nolimits[\xi_{k}(Z_{t_{k}})^{2}].$
Therefore
$W_{2}(\nu_{k+1},\nu_{k})\leq e^{\Gamma(T-t_{k+1})}\Bigl(\mathop{\mathbb{E}}\nolimits[\xi_{k}(Z_{t_{k}})^{2}]\Bigr)^{1/2}.$
Summing over $k=K,\dots,N-1$ yields
$W_{2}\!\bigl(\operatorname{Law}(Z_{t_{N}}),\,\mu_{K}P_{t_{s},T}\bigr)\leq\Delta_{\mathrm{late}}^{\mathrm{disc}}(s_{0}).$ (44)
switch-time discrepancy and late score forcing.
We separate the switch-time mismatch from the forcing accumulated after the switch:
$W_{2}(\mu_{K}P_{t_{s},T},p_{0})\leq W_{2}(\mu_{K}P_{t_{s},T},p_{s_{0}}P_{t_{s},T})+W_{2}(p_{s_{0}}P_{t_{s},T},p_{0}).$
For the first term, choose an optimal coupling $\pi\in\Pi(\mu_{K},p_{s_{0}})$, let $(Y_{t_{s}},\bar{Y}_{t_{s}})\sim\pi$, and let $Y,\bar{Y}$ solve the learned reverse SDE on $t_{s},T]$ driven by the same Brownian motion. Set
$D_{t}:=Y_{t}-\bar{Y}_{t}.$
Because the noise is synchronized,
$\dot{D}_{t}=\widehat{b}_{t}(Y_{t})-\widehat{b}_{t}(\bar{Y}_{t})\qquad\text{for a.e. }t\in[t_{s},T].$
By Lemma 1(iii),
$\langle D_{t},\widehat{b}_{t}(Y_{t})-\widehat{b}_{t}(\bar{Y}_{t})\rangle\leq\mathfrak{b}(T-t)\|D_{t}\|^{2}\qquad\text{for a.e. }t\in[t_{s},T].$
Applying Lemma 12 with $F_{t}:=0$ and $\beta(t):=\mathfrak{b}(T-t)$ yields
$\|D_{T}\|\leq e^{\int_{t_{s}}^{T}\mathfrak{b}(T-r)\,\mathop{}\!\mathrm{d}r}\|D_{t_{s}}\|=e^{\Gamma(s_{0})}\|D_{t_{s}}\|.$
Squaring, taking expectations, and using the optimality of $\pi$ give
$W_{2}(\mu_{K}P_{t_{s},T},p_{s_{0}}P_{t_{s},T})\leq e^{\Gamma(s_{0})}W_{2}(\mu_{K},p_{s_{0}})=e^{\Gamma(s_{0})}W_{2}(\operatorname{Law}(Z_{t_{s}}),p_{s_{0}}).$
It remains to control $W_{2}(p_{s_{0}}P_{t_{s},T},p_{0})$. Let $\widetilde{X}$ be the global ideal reverse solution from the theorem-level hypothesis. Because the learned reverse SDE is strongly well posed on $t_{s},T]$, on the same filtered space and with the same Brownian motion $\overline{B}$ that drives $\widetilde{X}$, there exists a unique strong solution $Y^{\star}$ of the learned reverse SDE on $t_{s},T]$ with
$Y^{\star}_{t_{s}}=\widetilde{X}_{t_{s}}.$
Since $\operatorname{Law}(\widetilde{X}_{t_{s}})=p_{s_{0}}$, the Markov property gives
$\operatorname{Law}(Y^{\star}_{T})=p_{s_{0}}P_{t_{s},T}.$
Moreover,
$\operatorname{Law}(\widetilde{X}_{r})=p_{T-r}\quad\text{for all }r\in[t_{s},T],\qquad\text{in particular}\qquad\operatorname{Law}(\widetilde{X}_{T})=p_{0}.$
Set
$D_{t}:=Y^{\star}_{t}-\widetilde{X}_{t}.$
Because the Brownian motions are identical,
$\dot{D}_{t}=\widehat{b}_{t}(Y^{\star}_{t})-\widehat{b}_{t}(\widetilde{X}_{t})+\delta_{t}(\widetilde{X}_{t})\qquad\text{for a.e. }t\in[t_{s},T].$
By Lemma 1(iii),
$\langle D_{t},\widehat{b}_{t}(Y^{\star}_{t})-\widehat{b}_{t}(\widetilde{X}_{t})\rangle\leq\mathfrak{b}(T-t)\|D_{t}\|^{2}\qquad\text{for a.e. }t\in[t_{s},T].$
Applying Lemma 12 with
$A_{t}:=\widehat{b}_{t}(Y^{\star}_{t})-\widehat{b}_{t}(\widetilde{X}_{t}),\qquad F_{t}:=\delta_{t}(\widetilde{X}_{t}),\qquad\beta(t):=\mathfrak{b}(T-t),$
and using $D_{t_{s}}=0$ gives
$\|D_{T}\|\leq\int_{t_{s}}^{T}e^{\int_{r}^{T}\mathfrak{b}(T-\tau)\,\mathop{}\!\mathrm{d}\tau}\|\delta_{r}(\widetilde{X}_{r})\|\,\mathop{}\!\mathrm{d}r=\int_{t_{s}}^{T}e^{\Gamma(T-r)}\|\delta_{r}(\widetilde{X}_{r})\|\,\mathop{}\!\mathrm{d}r.$
Taking the $L^{2}$-norm and using Minkowski’s integral inequality yields
$\bigl(\mathop{\mathbb{E}}\nolimits\|D_{T}\|^{2}\bigr)^{1/2}\leq\int_{t_{s}}^{T}e^{\Gamma(T-r)}\bigl(\mathop{\mathbb{E}}\nolimits\|\delta_{r}(\widetilde{X}_{r})\|^{2}\bigr)^{1/2}\,\mathop{}\!\mathrm{d}r.$
Also, since $\widetilde{X}_{r}\sim p_{T-r}$,
$\bigl(\mathop{\mathbb{E}}\nolimits\|\delta_{r}(\widetilde{X}_{r})\|^{2}\bigr)^{1/2}=g^{2}(T-r)\Bigl(\mathop{\mathbb{E}}\nolimits_{X\sim p_{T-r}}\|e_{T-r}(X)\|^{2}\Bigr)^{1/2}\leq g^{2}(T-r)\varepsilon(T-r)$
by Assumption 2. Therefore the coupling $(Y^{\star}_{T},\widetilde{X}_{T})$ gives
$W_{2}(p_{s_{0}}P_{t_{s},T},p_{0})\leq\bigl(\mathop{\mathbb{E}}\nolimits\|D_{T}\|^{2}\bigr)^{1/2}\leq\int_{t_{s}}^{T}e^{\Gamma(T-r)}g^{2}(T-r)\varepsilon(T-r)\,\mathop{}\!\mathrm{d}r.$
Changing variables s=Trs=T-r gives
$W_{2}(p_{s_{0}}P_{t_{s},T},p_{0})\leq\Delta_{\mathrm{late}}^{\mathrm{force}}(s_{0}).$ (45)
Combining the two bounds proves
$W_{2}(\mu_{K}P_{t_{s},T},p_{0})\leq e^{\Gamma(s_{0})}W_{2}(\operatorname{Law}(Z_{t_{s}}),p_{s_{0}})+\Delta_{\mathrm{late}}^{\mathrm{force}}(s_{0}).$ (46)
combine the two late pieces.
Combining (44) with (46) and the triangle inequality proves (43).
Appendix F Main theorem proof and Euclidean comparison
F.1 Proof of the main theorem and switch selection
Proof of Theorem 1.
Fix an admissible grid-aligned switch $s_{0}\in\mathcal{S}_{\mathrm{adm}}$, with $t_{K}=t_{s}:=T-s_{0}$. By Lemma 14,
$\Delta_{\varphi}^{\mathrm{sw}}(s_{0})\leq\overline{\Delta}_{\varphi}^{\mathrm{sw}}(s_{0}).$
By Lemma 15,
$W_{2}\!\bigl(\operatorname{Law}(Z_{t_{s}}),p_{s_{0}}\bigr)\leq C_{p}^{\mathrm{sw}}(s_{0})\,\bigl(\Delta_{\varphi}^{\mathrm{sw}}(s_{0})\bigr)^{\theta_{p}}\leq C_{p}^{\mathrm{sw}}(s_{0})\,\bigl(\overline{\Delta}_{\varphi}^{\mathrm{sw}}(s_{0})\bigr)^{\theta_{p}}.$
Finally, Proposition 4 yields
$W_{2}\!\bigl(\operatorname{Law}(Z_{t_{N}}),p_{0}\bigr)\leq\Delta_{\mathrm{late}}(s_{0})+e^{\Gamma(s_{0})}W_{2}\!\bigl(\operatorname{Law}(Z_{t_{s}}),p_{s_{0}}\bigr).$
Substituting the switch-time conversion estimate proves (26). ∎
Definition 7 (Scalar switch-selection objective).
For an admissible grid-aligned switch $s_{0}\in\mathcal{S}_{\mathrm{adm}}$, define
$\mathcal{B}_{p}(s_{0}):=\Delta_{\mathrm{late}}(s_{0})+e^{\Gamma(s_{0})}\,C_{p}^{\mathrm{sw}}(s_{0})\,\Bigl(\overline{\Delta}_{\varphi}^{\mathrm{sw}}(s_{0})\Bigr)^{\theta_{p}}.$ (47)
Corollary 2 (Discrete one-switch routing principle).
Define the admissible grid
$\mathcal{S}_{\mathrm{adm}}^{h}:=\mathcal{S}_{\mathrm{adm}}\cap\{T-t_{k}:\ k=1,\dots,N-1\}.$
Under the assumptions of Theorem 1, every switch $s_{0}\in\mathcal{S}_{\mathrm{adm}}^{h}$ satisfies
$W_{2}\!\bigl(\operatorname{Law}(Z_{t_{N}}),p_{0}\bigr)\leq\mathcal{B}_{p}(s_{0}).$
Hence, whenever $\mathcal{S}_{\mathrm{adm}}^{h}\neq\varnothing$, the one-switch routing problem on the numerical grid reduces to the scalar minimization
$\min_{s_{0}\in\mathcal{S}_{\mathrm{adm}}^{h}}\mathcal{B}_{p}(s_{0}).$
In particular, there exists $s_{0}^{\star}\in\mathcal{S}_{\mathrm{adm}}^{h}$ such that
$\mathcal{B}_{p}(s_{0}^{\star})=\min_{s_{0}\in\mathcal{S}_{\mathrm{adm}}^{h}}\mathcal{B}_{p}(s_{0}),\qquad W_{2}\!\bigl(\operatorname{Law}(Z_{t_{N}}),p_{0}\bigr)\leq\mathcal{B}_{p}(s_{0}^{\star}).$
Remark 8.
The substance of Corollary 2 is not the elementary fact that a minimum exists on a finite set. Its content is structural: Theorem 1 has already compressed an end-to-end geometric routing problem into the single scalar field
[…]
Admissibility is the geometric existence statement that the early window supports a positive far-field contraction reserve. Optimization is the quantitative routing statement, within the one-switch family studied here, that chooses where one should leave the first contracting geometry, pay the one-time conversion back to $W_{2}$, and expose the remaining discrepancy to the late Euclidean load. In this sense, the corollary turns a pathwise choice of interface between two geometries into a one-dimensional design problem on the admissible grid.
It also clarifies why an exact routed-versus-direct comparison is possible: for fixed $s_{0}$, the late-window contribution is the same in both bounds, so any strict improvement can only come from the early window. We do not claim that this scalar optimization is globally optimal among multi-switch or continuously varying metric strategies.
Proof of Corollary 2.
Theorem 1 together with Definition 7 yields
$W_{2}\!\bigl(\operatorname{Law}(Z_{t_{N}}),p_{0}\bigr)\leq\mathcal{B}_{p}(s_{0}),\qquad s_{0}\in\mathcal{S}_{\mathrm{adm}}^{h}.$
If $\mathcal{S}_{\mathrm{adm}}^{h}\neq\varnothing$, then $\mathcal{S}_{\mathrm{adm}}^{h}$ is finite, hence $\mathcal{B}_{p}$ attains a minimum on $\mathcal{S}_{\mathrm{adm}}^{h}$. Applying Theorem 1 at a minimizer gives the stated bound. ∎
F.2 Comparison with direct Euclidean propagation
Proposition 5.
Assume moreover that the ideal reverse SDE (12) admits a global strong solution $(\widetilde{X}_{t})_{0\leq t\leq T}$ with
$\operatorname{Law}(\widetilde{X}_{t})=p_{T-t},\qquad 0\leq t\leq T.$
Assume Assumptions 1, 2, and 3, that $\Gamma(T)<\infty$, and that the learned reverse SDE is strongly well posed on $[0,T]$. Assume moreover that the one-step numerical defects satisfy
$W_{2}\!\bigl(P_{k}(x,\cdot),\Psi_{k}(x,\cdot)\bigr)\leq\xi_{k}(x),\qquad x\in\mathbb{R}^{d},\quad k=0,\dots,N-1,$
and that the relevant second moments are finite. Then
$W_{2}\!\bigl(\operatorname{Law}(Z_{t_{N}}),p_{0}\bigr)\leq e^{\Gamma(T)}W_{2}(\widehat{p}_{T},p_{T})$ $+\sum_{k=0}^{N-1}e^{\Gamma(T-t_{k+1})}\Bigl(\mathop{\mathbb{E}}\nolimits[\xi_{k}(Z_{t_{k}})^{2}]\Bigr)^{1/2}$ (48)
$+\int_{0}^{T}e^{\Gamma(s)}\,g^{2}(s)\,\varepsilon(s)\,\mathop{}\!\mathrm{d}s.$
Proof.
Apply Proposition 4 with the candidate switch $s_{0}=T$. Then $t_{s}=0$, $K=0$, and $p_{s_{0}}=p_{T}$. The late-window bound becomes exactly (48). ∎
Corollary 3.
Assume the hypotheses of Theorem 1 and of Proposition 5. Fix an admissible grid-aligned switch $s_{0}\in\mathcal{S}_{\mathrm{adm}}$, with $t_{K}=t_{s}:=T-s_{0}$. Write
$d_{k}:=\Bigl(\mathop{\mathbb{E}}\nolimits[\xi_{k}(Z_{t_{k}})^{2}]\Bigr)^{1/2},\qquad\Lambda_{s_{0}}(r):=\int_{s_{0}}^{r}\mathfrak{b}(u)\,\mathop{}\!\mathrm{d}u,\qquad r\in[s_{0},T].$
Define the common late-window term
$\mathcal{L}(s_{0}):=\Delta_{\mathrm{late}}^{\mathrm{disc}}(s_{0})+\Delta_{\mathrm{late}}^{\mathrm{force}}(s_{0}),$
the routed early transport term
$\mathcal{R}_{\mathrm{sw}}(s_{0}):=C_{p}^{\mathrm{sw}}(s_{0})\Biggl[e^{-c(s_{0})t_{s}}W_{\varphi_{s_{0}}}(\widehat{p}_{T},p_{T})+\sum_{k=0}^{K-1}e^{-c(s_{0})(t_{s}-t_{k+1})}d_{k}+\int_{s_{0}}^{T}e^{-c(s_{0})(s-s_{0})}g^{2}(s)\varepsilon(s)\,\mathop{}\!\mathrm{d}s\Biggr]^{\theta_{p}},$
and the direct early transport term
$\mathcal{R}_{\mathrm{dir}}(s_{0}):=e^{\Lambda_{s_{0}}(T)}W_{2}(\widehat{p}_{T},p_{T})+\sum_{k=0}^{K-1}e^{\Lambda_{s_{0}}(T-t_{k+1})}d_{k}+\int_{s_{0}}^{T}e^{\Lambda_{s_{0}}(s)}g^{2}(s)\varepsilon(s)\,\mathop{}\!\mathrm{d}s.$
Then
$\mathcal{B}_{p}(s_{0})=\mathcal{L}(s_{0})+e^{\Gamma(s_{0})}\mathcal{R}_{\mathrm{sw}}(s_{0}),\qquad\mathcal{D}_{\mathrm{dir}}(s_{0})=\mathcal{L}(s_{0})+e^{\Gamma(s_{0})}\mathcal{R}_{\mathrm{dir}}(s_{0}),$
and hence
$W_{2}\!\bigl(\operatorname{Law}(Z_{t_{N}}),p_{0}\bigr)\leq\min\bigl\{\mathcal{B}_{p}(s_{0}),\,\mathcal{D}_{\mathrm{dir}}(s_{0})\bigr\}.$
Moreover,
$\mathcal{B}_{p}(s_{0})\leq\mathcal{D}_{\mathrm{dir}}(s_{0})\iff\mathcal{R}_{\mathrm{sw}}(s_{0})\leq\mathcal{R}_{\mathrm{dir}}(s_{0}),$
with strict inequality if and only if $\mathcal{R}_{\mathrm{sw}}(s_{0})<\mathcal{R}_{\mathrm{dir}}(s_{0})$.
Proof.
The identity
$\mathcal{B}_{p}(s_{0})=\mathcal{L}(s_{0})+e^{\Gamma(s_{0})}\mathcal{R}_{\mathrm{sw}}(s_{0})$
is just Definition 7. For the direct route, note that
$e^{\Gamma(T)}=e^{\Gamma(s_{0})}e^{\Lambda_{s_{0}}(T)},\qquad e^{\Gamma(T-t_{k+1})}=e^{\Gamma(s_{0})}e^{\Lambda_{s_{0}}(T-t_{k+1})},$
and
$e^{\Gamma(s)}=e^{\Gamma(s_{0})}e^{\Lambda_{s_{0}}(s)},\qquad s\in[s_{0},T].$
Substituting these identities gives
$\mathcal{D}_{\mathrm{dir}}(s_{0})=\mathcal{L}(s_{0})+e^{\Gamma(s_{0})}\mathcal{R}_{\mathrm{dir}}(s_{0}).$
The bound by $\min\{\mathcal{B}_{p}(s_{0}),\mathcal{D}_{\mathrm{dir}}(s_{0})\}$ combines Theorem 1 with Proposition 5, and the comparison criterion follows by subtracting the common term $\mathcal{L}(s_{0})$ and then the common factor $e^{\Gamma(s_{0})}$. ∎
Remark 9.
The comparison becomes transparent only after factoring out the common post-switch factor $e^{\Gamma(s_{0})}$. After that reduction, the two routes propagate the same three pre-switch inputs in opposite geometric directions. The direct route weights an input created at noise level $u\in[s_{0},T]$ by the residual Euclidean factor
$e^{\Lambda_{s_{0}}(u)},$
whereas the routed bound first damps the same input by
$e^{-c(s_{0})(u-s_{0})},$
and then pays a single return to $W_{2}$ through $C_{p}^{\mathrm{sw}}(s_{0})$ and $\theta_{p}$.
This identifies the genuine mismatch regime: radial contraction is already available, while direct Euclidean propagation on the same window is still net expansive. A convenient sufficient indicator is
$\underline{m}(s_{0})>0\qquad\text{and}\qquad\Lambda_{s_{0}}(T)=\int_{s_{0}}^{T}\mathfrak{b}(v)\,\mathop{}\!\mathrm{d}v>0.$
Then the routed bound damps the early inputs, whereas the direct bound amplifies them on net. The only competing loss is the one-time metric conversion. Accordingly, a strict improvement should be expected when the early window is long enough and sufficiently inside the admissible regime so that $c(s_{0})$ is appreciable and $C_{p}^{\mathrm{sw}}(s_{0})$ is not too large. By contrast, if the residual Euclidean load is already nonpositive, or if $s_{0}$ is chosen too close to the admissibility threshold, there is no reason to expect a systematic advantage from routing. This regime is made explicit in the variance-preserving example of Proposition˜8.
Appendix G Auxiliary comparison principles and sufficient conditions
Nothing in this section is needed for the proof of Theorem 1. It records auxiliary comparison principles and sufficient conditions: a relation between Assumption 3 and classical global score-regularity hypotheses, conservative admissibility certificates for switch selection, sufficient conditions for the early reflection-coupling hypothesis, and an explicit variance-preserving comparison result.
G.1 Relation of score-error monotonicity to global score regularity
Remark 10 (What Assumption 3 actually controls).
Assumption 3 is the theorem-level geometric input on the score error. Although it is stated in one-sided Lipschitz form, what the argument actually uses is more specific: for a pair $(x,y)$, the score-error contribution to pairwise distance evolution appears only through
$g^{2}(s)\,\langle e_{s}(x)-e_{s}(y),x-y\rangle.$
Accordingly, the radial lower profile $\underline{\kappa}_{s}$, the switch reserve $\mathfrak{m}(s)$, and the late Euclidean load $\mathfrak{b}(s)$ depend on $e_{s}$ only through the component of $e_{s}(x)-e_{s}(y)$ parallel to the separation vector xyx-y. Purely transverse or rotational components are invisible to this mechanism. In this sense Assumption 3 is a pairwise geometric monotonicity condition, closer to the one-sided dissipativity hypotheses standard in SDE analysis [Higham et al., 2002] than to a full global Lipschitz bound on the vector field. By contrast, global Lipschitz assumptions are stronger verification devices, and are also the natural hypotheses under which direct Euclidean $W_{2}$ arguments are typically closed [Gao et al., 2025, Gao and Zhu, 2024].
Lemma 16 (Sufficient condition for Assumption 3).
Suppose that for a.e. $s\in(0,T]$, one of the following holds:
  1. (a)
    the map $x\mapsto e_{s}(x)$ is $C^{1}$, and there exists a measurable $\bar{\ell}:(0,T]\to\mathbb{R}$ such that
    $\lambda_{\max}\!\Bigl(\tfrac{1}{2}\bigl(\nabla e_{s}(x)+\nabla e_{s}(x)^{\top}\bigr)\Bigr)\leq\bar{\ell}(s),\qquad x\in\mathbb{R}^{d};$
  2. (b)
    there exists a measurable $L_{e}:(0,T]\to[0,\infty)$ such that
    $\|e_{s}(x)-e_{s}(y)\|\leq L_{e}(s)\|x-y\|,\qquad x,y\in\mathbb{R}^{d};$
  3. (c)
    there exist measurable $L_{\theta},L_{\star}:(0,T]\to[0,\infty)$ such that
    $\|s_{\theta}(x,s)-s_{\theta}(y,s)\|\leq L_{\theta}(s)\|x-y\|,\qquad\|\nabla\log p_{s}(x)-\nabla\log p_{s}(y)\|\leq L_{\star}(s)\|x-y\|,\qquad x,y\in\mathbb{R}^{d}.$
Then Assumption 3 holds with
$\ell(s)=\bar{\ell}(s)\quad\text{in case {(a)}},\qquad\ell(s)=L_{e}(s)\quad\text{in case {(b)}},\qquad\ell(s)=L_{\theta}(s)+L_{\star}(s)\quad\text{in case {(c)}}.$
In particular, the symmetrized Jacobian bound, global Lipschitz control of $e_{s}$, and separate Lipschitz control of $s_{\theta}$ and $\nabla\log p_{s}$ all imply Assumption 3, but they are not necessary for it.
Proof.
Case (a) is the standard symmetrized-Jacobian criterion. Fix $s$ such that the assumption holds and let $x,y\in\mathbb{R}^{d}$. Writing $\gamma(\tau)=y+\tau(x-y)$, one has
$e_{s}(x)-e_{s}(y)=\int_{0}^{1}\nabla e_{s}(\gamma(\tau))(x-y)\,\mathop{}\!\mathrm{d}\tau.$
Taking the inner product with xyx-y, symmetrizing, and using the eigenvalue bound gives
$\langle e_{s}(x)-e_{s}(y),x-y\rangle\leq\bar{\ell}(s)\|x-y\|^{2}.$
For case (b), Cauchy–Schwarz gives
$\langle e_{s}(x)-e_{s}(y),x-y\rangle\leq\|e_{s}(x)-e_{s}(y)\|\,\|x-y\|\leq L_{e}(s)\|x-y\|^{2}.$
For case (c), since
$e_{s}(x)-e_{s}(y)=\bigl(s_{\theta}(x,s)-s_{\theta}(y,s)\bigr)-\bigl(\nabla\log p_{s}(x)-\nabla\log p_{s}(y)\bigr),$
the triangle inequality yields
$\|e_{s}(x)-e_{s}(y)\|\leq\bigl(L_{\theta}(s)+L_{\star}(s)\bigr)\|x-y\|,$
and case (b) applies. ∎
Remark 11 (Assumption 3 is strictly weaker than global Lipschitz).
The converse of Lemma 16 fails dramatically. Let
$e_{s}(x)=A_{s}x,\qquad A_{s}^{\top}=-A_{s},$
where $\|A_{s}\|_{\mathrm{op}}$ may be arbitrarily large. Then for every $x,y\in\mathbb{R}^{d}$,
$\langle e_{s}(x)-e_{s}(y),x-y\rangle=\langle A_{s}(x-y),x-y\rangle=0,$
so Assumption 3 holds with $\ell(s)=0$. However,
$\|e_{s}(x)-e_{s}(y)\|\leq\|A_{s}\|_{\mathrm{op}}\|x-y\|,$
and the global Lipschitz constant can be arbitrarily large. Thus Assumption 3 controls only the symmetric/expansive part of the score error, not the full vector increment. This is exactly why it is the right structural condition for pairwise distance propagation.
Remark 12 (Assumptions 2 and 3 are complementary).
Assumption 2 and Assumption 3 address different error channels. The former is an averaged forcing bound under $p_{s}$:
$\mathop{\mathbb{E}}\nolimits_{X\sim p_{s}}\|e_{s}(X)\|^{2}\leq\varepsilon^{2}(s).$
The latter is a pairwise geometric monotonicity bound:
$\langle e_{s}(x)-e_{s}(y),x-y\rangle\leq\ell(s)\|x-y\|^{2}.$
Neither replaces the other. A field may satisfy Assumption 3 with $\ell(s)=0$ and still have large $L^{2}(p_{s})$-size; the skew-symmetric example above already shows this as soon as $p_{s}$ has nontrivial second moment. Conversely, a field may have small averaged $L^{2}(p_{s})$-error while having a large positive one-sided slope on a region of small $p_{s}$-mass; such a field barely changes the forcing integral but can destroy pairwise contraction. So the two hypotheses are not redundant: Assumption 2 controls averaged forcing, while Assumption 3 controls pairwise expansion.
G.2 Conservative admissibility and switch-selection certificates
Proposition 6 (Proxy admissibility and proxy certificates).
Assume Assumptions 1,  2 and 3, and suppose that $\ell(s)\leq\bar{\ell}(s)$ and $\varepsilon(s)\leq\bar{\varepsilon}(s)$ for a.e. $s\in(0,T]$, where $\bar{\ell}:(0,T]\to\mathbb{R}$ and $\bar{\varepsilon}:(0,T]\to[0,\infty)$ are measurable. Define
$\mathfrak{m}_{\mathrm{pr}}(s):=g^{2}(s)\bigl(\alpha_{s}-\bar{\ell}(s)\bigr)-f(s),\qquad\mathfrak{b}_{\mathrm{pr}}(s):=f(s)+g^{2}(s)\bigl(M_{s}+\bar{\ell}(s)-\alpha_{s}\bigr).$
Then every grid-aligned $s_{0}$ with
$\underline{m}_{\mathrm{pr}}(s_{0}):=\inf_{u\in[s_{0},T]}\mathfrak{m}_{\mathrm{pr}}(u)>0$
is admissible, and the proof of Theorem 1 goes through with the conservative replacements
$\mathfrak{m},\mathfrak{b},\varepsilon\quad\rightsquigarrow\quad\mathfrak{m}_{\mathrm{pr}},\mathfrak{b}_{\mathrm{pr}},\bar{\varepsilon}.$
In particular, every such switch yields a certified proxy bound
$W_{2}\!\bigl(\operatorname{Law}(Z_{t_{N}}),p_{0}\bigr)\leq\mathcal{B}_{p}^{\mathrm{pr}}(s_{0}),$
where $\mathcal{B}_{p}^{\mathrm{pr}}(s_{0})$ is obtained from $\mathcal{B}_{p}(s_{0})$ by replacing the exact geometric inputs by their proxy versions.
Proof.
Introduce the proxy lower envelope
$\underline{\kappa}_{s}^{\mathrm{pr}}(r):=g^{2}(s)\Big(\alpha_{s}-\frac{1}{r}f_{M_{s}}(r)-\bar{\ell}(s)\Big)-f(s).$
Since $\ell(s)\leq\bar{\ell}(s)$, one has
$\underline{\kappa}_{s}(r)\geq\underline{\kappa}_{s}^{\mathrm{pr}}(r),\qquad r>0,\ s\in(0,T].$
Taking $r\downarrow 0$ and $r\to\infty$ shows
$\mathfrak{m}(s)\geq\mathfrak{m}_{\mathrm{pr}}(s),\qquad\mathfrak{b}(s)\leq\mathfrak{b}_{\mathrm{pr}}(s),\qquad s\in(0,T].$
Hence $\underline{m}_{\mathrm{pr}}(s_{0})>0$ implies $\underline{m}(s_{0})>0$, so every proxy-admissible switch is genuinely admissible.
The remaining claim follows by monotonicity of the constructions used in the proof: a smaller lower profile can only worsen the early contraction input, a larger forcing bound $\bar{\varepsilon}$ can only enlarge the forcing terms, and a larger Euclidean load $\mathfrak{b}_{\mathrm{pr}}$ can only enlarge the late propagation factor. Rerunning the early contraction, switch conversion, and late Euclidean propagation arguments with $\underline{\kappa}^{\mathrm{pr}},\mathfrak{m}_{\mathrm{pr}},\mathfrak{b}_{\mathrm{pr}},\bar{\varepsilon}$ therefore produces a certified upper bound $\mathcal{B}_{p}^{\mathrm{pr}}(s_{0})$. ∎
G.3 Sufficient conditions for the early reflection-coupling hypothesis
Proposition 7 (A convenient sufficient condition for the early coupling input).
Fix an admissible switch $s_{0}\in\mathcal{S}_{\mathrm{adm}}$ and write $t_{s}:=T-s_{0}$. Assume:
  1. (a)
    $g$ is continuous on $s_{0},T]$ and $\underline{g}(s_{0})>0$;
  2. (b)
    the map $(t,x)\mapsto\widehat{b}_{t}(x)$ is Borel measurable on $0,t_{s}]\times\mathbb{R}^{d}$;
  3. (c)
    there exists $L_{s_{0}}<\infty$ such that
    $\|\widehat{b}_{t}(x)-\widehat{b}_{t}(y)\|\leq L_{s_{0}}\|x-y\|,\qquad t\in[0,t_{s}],\ x,y\in\mathbb{R}^{d};$
  4. (d)
    there exists $B_{s_{0}}<\infty$ such that
    $\|\widehat{b}_{t}(x)\|\leq B_{s_{0}}(1+\|x\|),\qquad t\in[0,t_{s}],\ x\in\mathbb{R}^{d}.$
Then the learned reverse SDE (13) is strongly well posed on $0,t_{s}]$. Moreover, for every initial coupling $\pi\in\Pi(\mu,\nu)$ with $\mu,\nu\in\mathcal{P}_{1}(\mathbb{R}^{d})$ and every subinterval $u,v]\subseteq[0,t_{s}]$, there exists a coalescing reflection coupling in the sense of Definition 5; see Eberle [2016], Eberle et al. [2019].
Proof.
Assumptions (b)–(d) are the standard measurable, global-Lipschitz, and linear-growth hypotheses yielding strong existence and pathwise uniqueness for (13) on $0,t_{s}]$.
Fix a subinterval $u,v]\subseteq[0,t_{s}]$ and an initial coupling $\pi\in\Pi(\mu,\nu)$. Because the diffusion coefficient is the deterministic isotropic matrix
$g(T-t)I_{d},$
with $g$ continuous and bounded away from 0 on $[u,v]$ by (a), the standard reflection-coupling construction for isotropic diffusions with globally Lipschitz drift applies on $[u,v]$: one reflects the driving Brownian motion across the hyperplane orthogonal to the current separation vector up to the meeting time and switches to synchronous coupling afterwards. See Eberle [2016, Section 2] and Eberle et al. [2019, Section 2.1]. The present coefficient differs from the unit-diffusion case only by a deterministic scalar factor, so the same construction extends directly on each finite subinterval. Therefore, for every initial coupling $\pi$, there exists a coalescing reflection coupling in the sense of Definition 5. ∎
Corollary 4 (A score-model sufficient condition).
Fix an admissible switch $s_{0}\in\mathcal{S}_{\mathrm{adm}}$. Assume $g$ is continuous on $s_{0},T]$ with $\underline{g}(s_{0})>0$, that $f$ and $g$ are bounded on $s_{0},T]$, and that there exist constants $L_{\theta}(s_{0}),B_{\theta,0}(s_{0})<\infty$ such that
$\|s_{\theta}(x,s)-s_{\theta}(y,s)\|\leq L_{\theta}(s_{0})\|x-y\|,\qquad\|s_{\theta}(0,s)\|\leq B_{\theta,0}(s_{0}),\qquad s\in[s_{0},T],\ x,y\in\mathbb{R}^{d}.$
Then Proposition 7 applies. In particular, the coupling input required by Lemma 13 is automatic.
Proof.
For $t\in[0,t_{s}]$,
$\widehat{b}_{t}(x)-\widehat{b}_{t}(y)=f(T-t)(x-y)+g^{2}(T-t)\bigl(s_{\theta}(x,T-t)-s_{\theta}(y,T-t)\bigr),$
hence
$\|\widehat{b}_{t}(x)-\widehat{b}_{t}(y)\|\leq\Bigl(\|f\|_{L^{\infty}([s_{0},T])}+\|g\|_{L^{\infty}([s_{0},T])}^{2}L_{\theta}(s_{0})\Bigr)\|x-y\|.$
Likewise,
$\|\widehat{b}_{t}(x)\|\leq\|f\|_{L^{\infty}([s_{0},T])}\|x\|+\|g\|_{L^{\infty}([s_{0},T])}^{2}\Bigl(\|s_{\theta}(0,T-t)\|+L_{\theta}(s_{0})\|x\|\Bigr),$
so
$\|\widehat{b}_{t}(x)\|\leq B_{s_{0}}(1+\|x\|)$
for a suitable finite $B_{s_{0}}$. Thus all hypotheses of Proposition 7 are satisfied. ∎
G.4 Variance-preserving schedules: explicit routed-versus-direct certificates
The point of this example is not merely to compute proxy admissibility explicitly. It is to exhibit a concrete regime in which the phase-aware certificate is structurally sharper than the direct Euclidean certificate. For the variance-preserving schedule, the formulas become fully explicit. In the regime $\alpha<1$, Gaussian smoothing increases $\alpha_{s}$ monotonically, so both the opening of the admissible radial window and the routed-versus-direct comparison can be written in closed form.
Proposition 8 (VP proxy formulas and explicit routed-versus-direct certificates).
Consider the variance-preserving schedule
$f(s)\equiv\frac{\beta}{2},\qquad g(s)\equiv\sqrt{\beta},\qquad\beta>0,$
and assume the proxy bounds
$\ell(s)\leq\bar{\ell},\qquad\varepsilon(s)\leq\bar{\varepsilon},\qquad s\in(0,T],$
for constants $\bar{\ell}\in\mathbb{R}$ and $\bar{\varepsilon}\geq 0$. Let
$D(s):=\alpha+(1-\alpha)e^{-\beta s}.$
Assume moreover that the grid is uniform, $t_{k}=kh$, and that
$\Bigl(\mathop{\mathbb{E}}\nolimits[\xi_{k}(Z_{t_{k}})^{2}]\Bigr)^{1/2}\leq C_{\mathrm{sch}}h^{q},\qquad k=0,\dots,N-1.$
For a proxy-admissible grid-aligned switch $s_{0}$, let $t_{s}:=T-s_{0}=t_{K}$, $L:=T-s_{0}$, and let $\varphi_{s_{0}}^{\mathrm{pr}}$, $c_{\mathrm{pr}}(s_{0})$, and $C_{p,\mathrm{pr}}^{\mathrm{sw}}(s_{0})$ be the proxy switch metric, damping rate, and conversion constant furnished by Proposition 6 and (34)–(35) after replacing
$\mathfrak{m},\mathfrak{b},\varepsilon\quad\rightsquigarrow\quad\mathfrak{m}_{\mathrm{pr}},\mathfrak{b}_{\mathrm{pr}},\bar{\varepsilon}.$
Then the following hold.
  1. (i)
    The smoothed weak-log-concavity parameters are
    $a(s)=e^{-\beta s/2},\qquad\sigma^{2}(s)=1-e^{-\beta s},\qquad\alpha_{s}=\frac{\alpha}{D(s)},\qquad M_{s}=\frac{Me^{-\beta s}}{D(s)^{2}},$
    and the proxy reserve and load are
    $\mathfrak{m}_{\mathrm{pr}}(s)=\beta\Bigl(\frac{\alpha}{D(s)}-\bar{\ell}-\frac{1}{2}\Bigr),\qquad\mathfrak{b}_{\mathrm{pr}}(s)=\beta\Bigl(\frac{Me^{-\beta s}}{D(s)^{2}}+\bar{\ell}+\frac{1}{2}-\frac{\alpha}{D(s)}\Bigr).$
  2. (ii)
  3. (iii)
    Suppose $\alpha<1$ and $\bar{\ell}<\frac{1}{2}$. Define
    $s_{\mathrm{pr}}^{\star}:=\max\!\Biggl\{0,\;\frac{1}{\beta}\log\!\Bigl(\frac{(\bar{\ell}+\frac{1}{2})(1-\alpha)}{\alpha(\frac{1}{2}-\bar{\ell})}\Bigr)\Biggr\}.$
    Then every grid-aligned $s_{0}\in[s_{\mathrm{pr}}^{\star},T]$ is proxy-admissible, provided $s_{\mathrm{pr}}^{\star}\leq T$. Equivalently, proxy admissibility begins exactly when
    $\alpha_{s_{0}}>\bar{\ell}+\frac{1}{2}.$
    For later use, it is also convenient to record the primitive
    $\Gamma_{\mathrm{pr}}(s_{0}):=\int_{0}^{s_{0}}\mathfrak{b}_{\mathrm{pr}}(u)\,\mathop{}\!\mathrm{d}u=\beta\Bigl(\bar{\ell}+\frac{1}{2}\Bigr)s_{0}-\log\!\bigl(\alpha e^{\beta s_{0}}+1-\alpha\bigr)+\mathcal{M}_{\mathrm{VP}}(s_{0}),$
    where
    $\mathcal{M}_{\mathrm{VP}}(s_{0}):=\begin{cases}\displaystyle\frac{M}{1-\alpha}\Bigl(\frac{1}{D(s_{0})}-1\Bigr),&\alpha\neq 1,\\[10.00002pt] \displaystyle M(1-e^{-\beta s_{0}}),&\alpha=1.\end{cases}$
    The second branch is included only to record the closed-form primitive in the degenerate case $\alpha=1$.
  4. (iv)
    Define the common late-window proxy term
    $\Delta_{\mathrm{late}}^{\mathrm{pr}}(s_{0}):=C_{\mathrm{sch}}h^{q}\sum_{k=K}^{N-1}e^{\Gamma_{\mathrm{pr}}(T-t_{k+1})}+\beta\bar{\varepsilon}\int_{0}^{s_{0}}e^{\Gamma_{\mathrm{pr}}(u)}\,\mathop{}\!\mathrm{d}u,$
    and the routed early budget
    $\Xi_{\mathrm{sw}}^{\mathrm{pr}}(s_{0}):=e^{-c_{\mathrm{pr}}(s_{0})L}\,W_{\varphi_{s_{0}}^{\mathrm{pr}}}(\widehat{p}_{T},p_{T})+C_{\mathrm{sch}}h^{q}\,\frac{1-e^{-c_{\mathrm{pr}}(s_{0})L}}{1-e^{-c_{\mathrm{pr}}(s_{0})h}}+\frac{\beta\bar{\varepsilon}}{c_{\mathrm{pr}}(s_{0})}\Bigl(1-e^{-c_{\mathrm{pr}}(s_{0})L}\Bigr).$
    Then the phase-aware proxy certificate satisfies
    $W_{2}\!\bigl(\operatorname{Law}(Z_{t_{N}}),p_{0}\bigr)\leq\mathcal{B}_{p}^{\mathrm{pr}}(s_{0})\leq\Delta_{\mathrm{late}}^{\mathrm{pr}}(s_{0})+e^{\Gamma_{\mathrm{pr}}(s_{0})}\,C_{p,\mathrm{pr}}^{\mathrm{sw}}(s_{0})\,\Bigl(\Xi_{\mathrm{sw}}^{\mathrm{pr}}(s_{0})\Bigr)^{\theta_{p}}.$
  5. (v)
    Define the direct Euclidean proxy certificate by
    $\mathcal{D}_{\mathrm{dir,pr}}(s_{0}):=\Delta_{\mathrm{late}}^{\mathrm{pr}}(s_{0})+e^{\Gamma_{\mathrm{pr}}(T)}W_{2}(\widehat{p}_{T},p_{T})+C_{\mathrm{sch}}h^{q}\sum_{k=0}^{K-1}e^{\Gamma_{\mathrm{pr}}(T-t_{k+1})}+\beta\bar{\varepsilon}\int_{s_{0}}^{T}e^{\Gamma_{\mathrm{pr}}(u)}\,\mathop{}\!\mathrm{d}u.$
    If
    $\underline{\mathfrak{b}}_{\mathrm{pr}}(s_{0}):=\inf_{u\in[s_{0},T]}\mathfrak{b}_{\mathrm{pr}}(u)>0,$
    then
    $\mathcal{D}_{\mathrm{dir,pr}}(s_{0})-\Delta_{\mathrm{late}}^{\mathrm{pr}}(s_{0})\geq e^{\Gamma_{\mathrm{pr}}(s_{0})}\,\Xi_{\mathrm{dir}}^{\mathrm{pr}}(s_{0}),$
    where
    $\Xi_{\mathrm{dir}}^{\mathrm{pr}}(s_{0}):=e^{\underline{\mathfrak{b}}_{\mathrm{pr}}(s_{0})L}\,W_{2}(\widehat{p}_{T},p_{T})+C_{\mathrm{sch}}h^{q}\,\frac{e^{\underline{\mathfrak{b}}_{\mathrm{pr}}(s_{0})L}-1}{e^{\underline{\mathfrak{b}}_{\mathrm{pr}}(s_{0})h}-1}+\frac{\beta\bar{\varepsilon}}{\underline{\mathfrak{b}}_{\mathrm{pr}}(s_{0})}\Bigl(e^{\underline{\mathfrak{b}}_{\mathrm{pr}}(s_{0})L}-1\Bigr).$
    Consequently, whenever
    $C_{p,\mathrm{pr}}^{\mathrm{sw}}(s_{0})\,\Bigl(\Xi_{\mathrm{sw}}^{\mathrm{pr}}(s_{0})\Bigr)^{\theta_{p}}<\Xi_{\mathrm{dir}}^{\mathrm{pr}}(s_{0}),$
    one has the strict certificate-level improvement
    $\mathcal{B}_{p}^{\mathrm{pr}}(s_{0})<\mathcal{D}_{\mathrm{dir,pr}}(s_{0}).$
Proof.
Under the VP schedule,
$a(s)=\exp\!\Bigl(-\int_{0}^{s}\frac{\beta}{2}\,\mathop{}\!\mathrm{d}u\Bigr)=e^{-\beta s/2},$
and
$\sigma^{2}(s)=\int_{0}^{s}e^{-\beta(s-u)}\beta\,\mathop{}\!\mathrm{d}u=1-e^{-\beta s}.$
Substituting into (16) gives
$\alpha_{s}=\frac{\alpha}{\alpha+(1-\alpha)e^{-\beta s}}=\frac{\alpha}{D(s)},\qquad M_{s}=\frac{Me^{-\beta s}}{D(s)^{2}},$
which proves the formulas in (i). The expressions for $\mathfrak{m}_{\mathrm{pr}}$ and $\mathfrak{b}_{\mathrm{pr}}$ are the specialization of Proposition 6 to $f\equiv\beta/2$, $g^{2}\equiv\beta$, and the constant proxy bound $\bar{\ell}$.
Assume now $\alpha<1$. Then $s\mapsto\alpha_{s}$ is increasing, hence so is $s\mapsto\mathfrak{m}_{\mathrm{pr}}(s)$. Therefore
$\underline{m}_{\mathrm{pr}}(s_{0})>0\iff\mathfrak{m}_{\mathrm{pr}}(s_{0})>0\iff\frac{\alpha}{D(s_{0})}>\bar{\ell}+\frac{1}{2}.$
Solving this inequality gives exactly $s_{0}\geq s_{\mathrm{pr}}^{\star}$, proving the admissibility-threshold claim in (ii).
Independently of that monotonicity argument, the primitive $\Gamma_{\mathrm{pr}}$ is obtained from
$\Gamma_{\mathrm{pr}}(s_{0})=\beta\Bigl(\bar{\ell}+\frac{1}{2}\Bigr)s_{0}+\beta\int_{0}^{s_{0}}\frac{Me^{-\beta u}}{D(u)^{2}}\,\mathop{}\!\mathrm{d}u-\beta\int_{0}^{s_{0}}\frac{\alpha}{D(u)}\,\mathop{}\!\mathrm{d}u.$
The two integrals are explicit:
$\beta\int_{0}^{s_{0}}\frac{\alpha}{D(u)}\,\mathop{}\!\mathrm{d}u=\log\!\bigl(\alpha e^{\beta s_{0}}+1-\alpha\bigr),$
and
$\beta\int_{0}^{s_{0}}\frac{Me^{-\beta u}}{D(u)^{2}}\,\mathop{}\!\mathrm{d}u=\mathcal{M}_{\mathrm{VP}}(s_{0}).$
This gives the stated closed form, with the $\alpha=1$ branch obtained by direct integration.
For (iii), the late-window proxy term is just the split late contribution under the same conservative replacements. On the early window,
$\sum_{k=0}^{K-1}e^{-c_{\mathrm{pr}}(s_{0})(t_{s}-t_{k+1})}\Bigl(\mathop{\mathbb{E}}\nolimits[\xi_{k}(Z_{t_{k}})^{2}]\Bigr)^{1/2}\leq C_{\mathrm{sch}}h^{q}\sum_{k=0}^{K-1}e^{-c_{\mathrm{pr}}(s_{0})(t_{s}-t_{k+1})}.$
Since $t_{s}=t_{K}=Kh$,
$t_{s}-t_{k+1}=(K-k-1)h,$
and therefore
$\sum_{k=0}^{K-1}e^{-c_{\mathrm{pr}}(s_{0})(t_{s}-t_{k+1})}=\sum_{j=0}^{K-1}e^{-c_{\mathrm{pr}}(s_{0})jh}=\frac{1-e^{-c_{\mathrm{pr}}(s_{0})L}}{1-e^{-c_{\mathrm{pr}}(s_{0})h}}.$
Because $\bar{\varepsilon}$ is constant,
$\int_{s_{0}}^{T}e^{-c_{\mathrm{pr}}(s_{0})(u-s_{0})}\,\beta\bar{\varepsilon}\,\mathop{}\!\mathrm{d}u=\frac{\beta\bar{\varepsilon}}{c_{\mathrm{pr}}(s_{0})}\Bigl(1-e^{-c_{\mathrm{pr}}(s_{0})L}\Bigr).$
Substituting these estimates into Proposition 6 proves the routed proxy certificate.
For (iv), $\mathcal{D}_{\mathrm{dir,pr}}(s_{0})$ is the direct full-horizon Euclidean certificate obtained by making the same conservative replacements in Proposition 5 and splitting at the same index $K$. If $\underline{\mathfrak{b}}_{\mathrm{pr}}(s_{0})>0$, then for every $r\in[s_{0},T]$,
$\Gamma_{\mathrm{pr}}(r)-\Gamma_{\mathrm{pr}}(s_{0})=\int_{s_{0}}^{r}\mathfrak{b}_{\mathrm{pr}}(u)\,\mathop{}\!\mathrm{d}u\geq\underline{\mathfrak{b}}_{\mathrm{pr}}(s_{0})\,(r-s_{0}).$
Applying this to $r=T$, $r=T-t_{k+1}$, and inside the forcing integral gives
$e^{\Gamma_{\mathrm{pr}}(T)}\geq e^{\Gamma_{\mathrm{pr}}(s_{0})}e^{\underline{\mathfrak{b}}_{\mathrm{pr}}(s_{0})L},$
$\sum_{k=0}^{K-1}e^{\Gamma_{\mathrm{pr}}(T-t_{k+1})}\geq e^{\Gamma_{\mathrm{pr}}(s_{0})}\sum_{j=0}^{K-1}e^{\underline{\mathfrak{b}}_{\mathrm{pr}}(s_{0})jh}=e^{\Gamma_{\mathrm{pr}}(s_{0})}\frac{e^{\underline{\mathfrak{b}}_{\mathrm{pr}}(s_{0})L}-1}{e^{\underline{\mathfrak{b}}_{\mathrm{pr}}(s_{0})h}-1},$
and
$\int_{s_{0}}^{T}e^{\Gamma_{\mathrm{pr}}(u)}\,\mathop{}\!\mathrm{d}u\geq e^{\Gamma_{\mathrm{pr}}(s_{0})}\int_{0}^{L}e^{\underline{\mathfrak{b}}_{\mathrm{pr}}(s_{0})v}\,\mathop{}\!\mathrm{d}v=e^{\Gamma_{\mathrm{pr}}(s_{0})}\frac{e^{\underline{\mathfrak{b}}_{\mathrm{pr}}(s_{0})L}-1}{\underline{\mathfrak{b}}_{\mathrm{pr}}(s_{0})}.$
This proves the lower bound on $\mathcal{D}_{\mathrm{dir,pr}}(s_{0})-\Delta_{\mathrm{late}}^{\mathrm{pr}}(s_{0})$, and the strict comparison criterion follows immediately. ∎
Remark 13 (Long mismatch windows).
This example makes the advantage of routing completely explicit. If $c_{\mathrm{pr}}(s_{0})L\gg 1$, then the routed early budget has already saturated: the terminal mismatch carries the factor $e^{-c_{\mathrm{pr}}(s_{0})L}$, the pre-switch discretization term is of order $C_{\mathrm{sch}}h^{q-1}/c_{\mathrm{pr}}(s_{0})$, and the forcing term is of order $\beta\bar{\varepsilon}/c_{\mathrm{pr}}(s_{0})$. By contrast, if $\underline{\mathfrak{b}}_{\mathrm{pr}}(s_{0})>0$, then the direct Euclidean certificate transports the same three inputs with positive exponential weights over the same window. So on a genuine mismatch window—radial damping already active, residual Euclidean load still positive—the routed certificate can be strictly sharper, provided the one-time conversion cost $C_{p,\mathrm{pr}}^{\mathrm{sw}}(s_{0})$ is not too large.